### v1

# End-to-End 4D Heart Mesh Recovery Across Full-Stack and Sparse Cardiac MRI

Abstract
Reconstructing cardiac motion from cine CMR sequences is critical for diagnosis, prediction, and intervention. Existing methods rely on complete CMR stacks to infer full heart motion, limiting their utility in intra-procedural scenarios where only sparse observations are available. We present TetHeart, the first end-to-end framework that unifies full 4D multi-structure heart mesh recovery from both offline full-stack acquisitions and intra-procedural sparse-slice observations. Our method leverages deep deformable tetrahedra, an explicit–implicit hybrid representation, to capture shape and motion in a coherent space shared across cardiac structures. It is initialized from high-quality pre-procedural or offline-acquired full stacks to build detailed, patient-specific heart meshes, which can then be updated using whatever slices are available, from full stacks down to a single slice. We further incorporate several key innovations: (i) an attentive mechanism for slice-adaptive 2D–3D feature assembly that dynamically integrates information from arbitrary numbers of slices at any position, combined with a distillation strategy from full-slice to sparse-slice settings to ensure accurate reconstruction under extreme sparsity; and (ii) a two-stage weakly supervised motion learning scheme requiring only keyframe (e.g., ED and ES) annotations. Trained and validated on three large public datasets and externally evaluated zero-shot on additional private interventional and public CMR datasets, TetHeart achieves state-of-the-art accuracy and strong generalization in both pre- and intra-procedural settings.
{IEEEkeywords}
cardiac magnetic resonance (CMR), deformable tetrahedra, explicit–implicit hybrid representation, shape and motion modeling, sparse observation.
1 Introduction
The heart is one of the most vital organs in the human body [1], and accurate modeling of its shape and motion is central to numerous medical applications ranging from diagnosis to therapy planning [2, 3, 4]. Cardiac magnetic resonance (CMR) offers rich temporal imaging of cardiac anatomy and has become an indispensable modality for both shape [5, 6, 7, 8, 9] and motion [10, 11, 12, 13, 14] modeling. However, to this day, such modeling remains confined to offline analysis and pre-operative planning. Existing techniques [5, 6, 7, 8, 9, 10, 11, 12, 13, 14] cannot be applied during an intervention, as they require full-stack CMR data that cannot be acquired in real time. In the intervention room, only a handful of real-time slices can be captured, usually at lower resolution [15, 16, 17], a limitation also confirmed in routine practice at our collaborating university hospital. This scenario motivates us to move beyond the current state of the art and recover full 3D shape and motion from sparse and partial observations. This will enable real-time tracking of patient-specific heart dynamics, with a view to delivering guidance to interventional teams.
To this end, we introduce TetHeart, a unified framework for reconstructing the full 4D motion of multiple cardiac structures from either complete CMR stacks or intra-procedural sparse data. As in Fig. 1, TetHeart first leverages high-quality pre-operative or offline-acquired CMR stacks to build a detailed, patient-specific heart mesh, and then updates it in an end-to-end manner using any available slices, from a complete stack down to a single slice. To make this possible, TetHeart builds on deep deformable tetrahedra [18], an explicit–implicit hybrid representation that unites the optimization efficiency of signed distance functions with the geometric manipulability of meshes, forming the foundation for jointly modeling shape and motion in a coherent space shared across structures. On top of this, our first core contribution is an attentive mechanism for slice-adaptive 2D-3D feature assembly, which enables the model to handle sparse input by dynamically integrating information from arbitrary slices at arbitrary positions. Additionally, we introduce a distillation strategy from full-slice to sparse-slice settings, ensuring accurate reconstruction under extreme input sparsity. Our second contribution is a two-stage weakly supervised motion learning scheme that requires only keyframe (e.g., ED and ES) annotations, making it feasible to train TetHeart across diverse clinical environments. The core idea is to first train a static shape reconstruction branch on labeled frames, and then repurpose it to initialize shapes for unlabeled frames, from which motion dynamics are recovered on the labeled frames to provide supervision.
Figure 1: TetHeart reconstructs cardiac motion from observed 2D slices. At $t=0$, a 3D CMR stack $\mathbf{I}^{0}$ is processed by the AFA module with self-attention to extract a volumetric static feature $V_{\text{static}}^{0}$, which is then used to generate the initial tetrahedra $\mathcal{G}^{0}$. For each subsequent time step $t$, the observed 2D slices $\mathbf{O}^{t}$ are encoded via the AFA module with cross-attention to obtain a volumetric motion feature $V_{\text{motion}}^{t}$, which is then used to recover motion by predicting offsets that deform $\mathcal{G}^{0}$. Weights are shared between the static reconstruction branch and dynamic deformation branch.
By representing the heart in a shared deformable space, TetHeart preserves spatial alignment between structures while adapting to varying sparsity levels. This offline/online capability enables comprehensive, patient-specific motion tracking across the entire clinical workflow: not only achieving state-of-the-art accuracy for offline use with full-stack CMR, but also, more importantly, unlocking online use in intra-procedural scenarios where only sparse observations are available.
We conducted extensive experiments to train and validate our model on three large public datasets: ACDC [19], M&Ms [20], and M&Ms-2 [21]. Remarkably, Remarkably, with a single set of weights learned from only end-diastolic (ED) and end-systolic (ES) annotations, our method generalizes seamlessly across both online few-slice and offline full-slice settings, consistently outperforming existing approaches not only on geometric metrics (Chamfer Distance and Dice) but also on clinically relevant indexes related to ejection fraction. We further performed external evaluations without retraining to assess generalizability. First, we tested the method on a privately collected dataset tailored for the online interventional MRI scenario at our collaborating cardiac CMR center, demonstrating its potential for deployment in real-world clinical environments. In addition, we evaluated the model on the public 4DM dataset [14], which provides complete temporal annotations of the myocardium, thereby demonstrating that the proposed weakly supervised motion learning scheme enables accurate reconstruction across the entire cardiac cycle. Together, these results highlight the effectiveness, robustness, and clinical promise of TetHeart. We will make our code public.
2 Related Work
2.1 Static Cardiac Shape Modeling
Reconstructing the heart from 3D medical images holds significant clinical value. However, cardiac MR (CMR) images often exhibit low through-plane resolution, which makes accurate reconstruction a challenge. Over the years, many approaches to addressing it have been proposed [22, 5, 23, 24, 25, 26, 27, 28, 29, 9] . They can be categorized into traditional optimization-based techniques and deep learning-based methods.
2.1.1 Optimization-Based Approaches
They usually rely on a two-stage pipeline. Contours are first extracted from the CMR images—either manually or using a segmentation algorithm—and then a template mesh is optimized to fit these contours. For example, the approach of [23] starts from a predefined tubular initial mesh that is deformed to minimize a contour-matching loss. More recently, a 3D active shape model (3D-ASM) and an image intensity model was used in [29] to align initial meshes with ground-truth contours. While these approaches can produce good results, they also have drawbacks. As they rely on test-time optimization, the reconstruction process is often time-consuming, taking from tens of seconds to several minutes per mesh. When applied to a full image sequence, this may extend to over an hour, making them impractical for real-time applications. Moreover, these methods’ performance has only been evaluated using ground-truth contours, which require labor-intensive manual annotation. What their performance is when using predicted contours instead remains an open question.
2.1.2 Deep Learning-Based Approaches
Deep learning-based methods leverage prior knowledge learned from large datasets to accelerate mesh reconstruction. For instance, in [28], sets of contours are treated as sparse point cloud and decoded into 3D meshes using point- and graph-convolutions. Similarly, in [26], contour points are embedded into a 3D volume and a hybrid architecture made of 3D CNNs and GCNs is used to progressively deform a template mesh. These methods increase speed but also suffer from several limitations. Because meshes are reconstructed solely from extracted contours or points, the rich appearance information contained in cine images is not fully exploited. Furthermore, as contour extraction is not necessarily perfect, such two-states approaches can result in error accumulation [26].
2.2 Dynamic Cardiac Motion Modeling
Arguably, when reconstructing deforming hearts from sequences of CMR scans, one could simply run a static reconstruction algorithm on each individual frame. However, this would fail to exploit temporal consistency and make it difficult to establish inter-frame correspondences. Thus a number of motion-based approaches that compute the deformation from frame to frame have been proposed.
2.2.1 Motion from Full-Slice Offline MRI Sequences
When full CMR stacks are available, recent deep learning-based approaches [11, 12, 13, 14, 6] can be used. MulViMotion [12] trains 3D CNNs to estimate left ventricular myocardial motion by predicting voxel-based deformation fields from CMR data, and then building 3D meshes by warping segmentation masks. DeepMesh [13] decouples mesh reconstruction from motion estimation, and uses 3D CNNs and bilinear interpolation to predict per-vertex deformation. 4DMR [14] models the myocardium using implicit representation. Decoupled shape and motion latent codes are optimized during test time to predict an initial shape and the backward motion fields. We use MulViMotion, DeepMesh and 4DMR as baselines because they are among the best currently available.
Unfortunately, these approaches also suffer from several limitations. First, some of them focus only on myocardium, neglecting other cardiac structures, preventing comprehensive heart assessment. Second, optimization-based methods like 4DMR are time-consuming at inference time. And, finally and most importantly, they require full scans that cannot be acquired at sufficient frame rate, which precludes intra-procedural use as discussed below.
2.2.2 Motion from Sparse Online MRI Sequences
CMR imaging has a reputation for being a ”slow” modality. Acquiring a full MRI stack involves sequential slice acquisition over a 4–6 heartbeat breath-hold, which can take several minutes [16, 17]. While this is feasible for pre-interventional planning, it is not suitable for intra-procedural settings.
Recently, real-time MRI (RT-MRI) has gained favor due to its ability to capture dynamic processes without the need for gating, synchronization, or repetition [16]. Its applications include complex electrophysiology procedures [30], congenital heart disease [31], and other cardiac interventions [15, 32]. RT-MRI provides excellent soft-tissue contrast for assessing interventions like ablations [33, 17] and allows for continuous visualization and tracking of devices, and it is free of ionizing radiation [30, 34].
Nevertheless, although RT-MRI enables rapid and continuous image acquisition, it still involves a fundamental tradeoff between spatial resolution, temporal resolution, artifacts and reconstruction latency [16]. Therefore, achieving a sufficient temporal resolution (e.g., 20–25 Hz) to be used during invasive procedures, can only be done at the cost of acquiring a limited number of 2D slices at lower spatial resolution [15, 32, 34, 16, 17]. We are not aware of any current method that can reconstruct 3D cardiac motion using this kind of data. Yet, delivering a high-quality, dynamic 3D heart model to the intervention room based on real-time observations is crucial for guiding high-precision minimally invasive cardiac procedures [35, 36, 15] and for real-time localization of devices within the heart [30, 37]. This is the application that motivated us to begin this study.
3 Method
In this section, we introduce TetHeart, our framework for accurate 4D mesh reconstruction from either full CMR scans or sparse sets of 2D slices, as depicted by Fig 1. We first introduce our heart presentation that is based on deep deformable tetrahedra [18], an explicit–implicit hybrid representation that unites the optimization efficiency of signed distance functions with the geometric manipulability of meshes. We then introduce our Attentive 2D–3D Feature Assembler (AFA) module, which enables effective feature encoding from 2D slices with varying numbers and locations. Finally, we present our weakly supervised training scheme that makes it possible to train given only a limited number of annotations.
3.1 Formalization
To model deforming hearts, we use the deep deformable tetrahedra [18] formalism. We discretize the 3D space using a deformable tetrahedral grid, in which each vertex possesses an SDF value, denoted as $\mathcal{G}=(\{\mathbf{v}_{i}\},\{s_{i}\},\mathcal{T})$, where $\mathbf{v}_{i}$ and $s_{i}$ are the vertices and their corresponding SDF values in $\mathcal{T}$, the set of all tetrahedrons. The explicit surface can then be extracted using the differentiable marching tetrahedra algorithm [38].
Given an initial 3D heart model $\mathcal{G}^{0}$ reconstructed from the end-diastolic (ED) frame $\mathbf{I}^{0}\in\mathbb{R}^{D\times H\times W}$, our goal is to estimate a deformed model $\mathcal{G}^{t}$ at each time instant $t$, based on subsequent observations $\mathbf{O}^{t}$ that reflect the ongoing cardiac motion. We do this by updating the vertices of $\mathcal{G}^{0}$. We write
$\mathcal{G}^{t}=(\{\mathbf{v}_{i}^{t}\},\{s_{i}\},\mathcal{T})\;,\mathbf{v}_{i}^{t}=\mathcal{D}(\mathbf{v}_{i}^{0},\mathbf{O}^{t})\;,$ (1)
where $\mathbf{v}_{i}^{0}$ is a grid vertex from $\mathcal{G}^{0}$, $\mathbf{v}_{i}^{t}$ its counterpart after deformation. $\mathcal{D}$ is the deformation model which we instantiate in the following sections and makes it possible to recover deformations across the whole sequence.
The observations $\mathbf{O}^{t}$ can be obtained in two manners.
  1. 1.
    Online. For intra-procedural use, $\mathbf{I}^{0}\in\mathbb{R}^{D\times H\times W}$ is a full-slice image acquired before the intervention, taking as much time as necessary. However, it has to be possible to acquire $\mathbf{O}^{t}$ at sufficient frame rate and, consequently, it can only comprise a small set—from 1 to 3—of 2D slices at lower spatial resolution and at arbitrary locations [15].
  2. 2.
    Offline. Full-slice CMR images are available both initially and at time $t$, which is acquired from a slow slice-by-slice scanning process. In this case, $\mathbf{O}^{t}$ and $\mathbf{I}^{0}$ are of the same dimension $\mathbb{R}^{D\times H\times W}$.
In the following sections, we first introduce the AFA module and then describe how it is incorporated in our approach to generate the initial shape model from $\mathbf{I}^{0}$ and then deforming it given $\mathbf{O}^{t}$. Whether it is a full scan or a sparse set of slices, we use the same approach in both cases.
3.2 Attentive 2D-3D Feature Assembler
Because the observations can be either a full CMR stack, a sparse set of slices, or anything in between, we need a mechanism to extract useful features in all cases. To this end, we developed the Attentive 2D-3D Feature Assembler (AFA) module that relies on an attention mechanism that takes $\mathbf{I}^{0}$ and $\mathbf{O}^{t}$ as input and produces features.
Let us write $\mathbf{O}^{t}$ as $\{\mathbf{I}^{t}_{s}\in\mathbb{R}^{H\times W}|s=1,\ldots,S\}$, a collection of 2D slices where $S$ is the total number of slices. With this notation, $\mathbf{I}^{0}$ can be written as $\{\mathbf{I}^{0}_{d}\in\mathbb{R}^{H\times W}|d=1,\ldots,D\}$. A full-slice observation is simply one where $S=D$.
In practice, we first perform 2D convolutions on each individual slice to create a collection of 2D feature maps $F^{0}_{2d}$ and $F^{t}_{2d}$ from $\mathbf{I}^{0}$ and $\mathbf{O}^{t}$. Given that $\mathbf{O}^{t}$ is not necessarily in the same coordinate system as $\mathbf{I}^{0}$, we use the metadata associated to these slices in the DICOM header to position them spatially in the $\mathbf{I}^{0}$ volume. This links each 2D location in the feature maps to a precise 3D location. In the remainder of this section, we first discuss how we use these 2D feature maps to compute volumetric motion features from $F^{0}_{2d}$ and $F^{t}_{2d}$, which are obtained from slices acquired at different times and can be used to estimate deformations. We then describe how we derive static features from a dense set of slices all acquired at the same time, which can be used to reconstruct the initial static model.
Motion Features. A naive way to use an attention mechanism to construct motion features would be to treat the $F^{0}_{2d}$ as the query and the $F^{t}_{2d}$ as both keys and values. However, this would be inefficient because global attention aims at capturing long-range dependencies, which is not needed to model the inherently local nature of anatomical motion. Also, this requires computing full pairwise relationships between all query and key positions, which is computationally expensive when using many slices and high-resolution feature maps.
To avoid this, we take our inspiration from the Swin Transformer [39] and use instead a variant of attention that preserves locality and significantly reducing computational complexity. For each voxel in $F^{0}_{2d}$, we first identify $k_{1}$ spatially nearest 2D slices, or all slices if there are only $k_{1}$ or less. Within each selected slice, we then retrieve the features at the $k_{2}$ spatially closest positions to serve as keys and values. Attention is subsequently computed only between the query and its corresponding set of localized keys and values. This can be written as
$V^{t}_{\rm motion}=AFA(F^{0}_{2d},F^{t}_{2d})=\text{MHAttention}(Q,K,V)\;,$ (2)
$\text{where }Q=F^{0}_{2d},~K=V=\text{NN-Select}_{F^{0}_{2d},k_{1},k_{2}}(F^{t}_{2d})\;,$
where NN-Select denotes the nearest slice and position selecting operation. MHAttention denotes a Multi-Head Attention [40] operation. A learnable position embedding is used to encode the spatial information into the calculation of attention. In practice, we set $k_{1}=3$ and $k_{2}=9$. This configuration ensures that the computational complexity of our attention mechanism is approximately equivalent to that of a 3D convolution with a kernel size of 3. As a result, the AFA module achieves high computational efficiency while retaining the flexibility and adaptability of attention-based feature aggregation.
While our AFA module effectively encodes features from varying 2D slice configurations, a persistent challenge remains in the online scenario, where the limited number of slices can hinder accurate 3D motion reconstruction. To alleviate this issue, we introduce a distillation loss that encourages the model to efficiently transfer knowledge from full-slice inputs to partial-slice inputs during training. In short, we predict two versions of motion features from full-slice observation and its randomly sampled subset. The distillation loss encourages the encoded feature from partial slices to be close to the feature from full-slice input. We formulate the loss we minimize to achieve this in Section 3.5.
Static Features. The motion features described above relate $\mathbf{I}^{0}$ to $\mathbf{O}^{t}$. The same mechanism can also be used to relate $\mathbf{I}^{0}$ to itself and produce features that can be used for static reconstruction. This can be similarly written as
$V^{0}_{\rm static}=AFA(F^{0}_{2d},F^{0}_{2d})=\text{MHAttention}(Q,K,V)\;,$ (3)
$\text{where }Q=F^{0}_{2d},~K=V=\text{NN-Select}_{F^{0}_{2d},k_{1},k_{2}}(F^{0}_{2d})\;.$
Generating Final Features. After obtaining the volumetric static and motion features $V^{0}_{\rm static}$ and $V^{t}_{\rm motion}$, we use a U-Net-like [41, 42] architecture for further encoding. Just as standard U-Net architecture, it also consists of a downsampling and an upsampling stream. Features from the last level of the upsample stream are taken as the final features, which we denoted as $F^{0}_{\rm static}$ and $F^{t}_{\rm motion}$, respectively.
Note that the initial 2D encoding block and the subsequent 3D convolutions are shared for both static and motion feature extraction. This parameter sharing strategy enabled by our unified design allows motion prediction to benefit from the rich 3D spatial information learned in the static reconstruction task. Consequently, it facilitates more accurate motion estimation—an advantage overlooked by some previous works [13, 14]. As demonstrated in the experiments, this architecture leads to superior motion reconstruction performance.
3.3 Initial Heart Model Reconstruction
Given the $F^{0}_{\rm static}$ features computed as discussed above from a full-slice CMR image $\mathbf{I}^{0}$, our first task is to reconstruct the initial tetrahedral mesh $\mathcal{G}^{0}$. To this end, we start from an initial tetrahedral grid $\mathcal{G}^{init}$ obtained by uniformly sampling a unit cube. We then trilinearly interpolate $F^{0}_{\rm static}$ at the vertices of $\mathcal{G}^{init}$. An SDF value and an offset for each vertex is then predicted using a GCN to create $\mathcal{G}^{0}$. We write
$\mathcal{G}^{0}$ $=(\{\mathbf{v}_{i}^{\prime}\},\{s_{i}\},T),$
$\mathbf{v}_{i}^{\prime}$ $=\mathbf{v}_{i}+\Delta\mathbf{v}_{i},$ (4)
$(\Delta\mathbf{v}_{i},s_{i})$ $=\operatorname{GCN}([F^{0}_{\rm static}(\mathbf{v}_{i}),\mathbf{v}_{i}])\;,$
where $\mathbf{v}_{i}$ is a grid vertex of $\mathcal{G}^{init}$ and $\mathbf{v}_{i}^{\prime}$ the same grid vertex after deformation. $F^{0}_{\rm static}(\mathbf{v})$ denotes feature extraction at $\mathbf{v}$ from $F^{0}_{\rm static}$ with trilinear interpolation, and $\cdot$$\cdot$] denotes concatenation. To enhance the network’s spatial sensitivity, the vertex location is appended to the feature vector before it is passed to the GCN. For datasets with $C$ classes, we simply set the output channel of the GCN layer to $C\times 4$, 3 for deformation, and 1 for SDF value. This yields a separate tetrahedral mesh for each class. Following prior works [43, 44, 45, 46], our network also has an auxiliary segmentation prediction branch during training time.
3.4 Motion Recovery
The static reconstruction scheme described above delivers comparable or even better reconstruction performance to that of approaches based on pure 3D U-Net architectures [43, 44, 45, 46], with the added benefit that it also allows the independent encoding of individual 2D slices. We now exploit this mechanism to compute deformations for the initial static shape from sets of slices acquired later. Given $\mathcal{G}^{0}$, $F^{0}_{\rm static}$, and $F^{t}_{\rm motion}$ computed as described above, we use a combination of GCN and GRU [47] layers to effectively aggregate spatial information and instantiate the deformation model $\cal{D}$ of Eq. 1. We write
$F^{cat}(\mathbf{v}_{i}^{(s)})$ $=[F^{0}_{\text{static}}(\mathbf{v}_{i}^{(s)}),F^{t}_{\text{motion}}(\mathbf{v}_{i}^{(s)})]\;,$ (5)
$F^{gcn}(\mathbf{v}_{i}^{(s)})$ $=\operatorname{GCN}(F^{cat}(\mathbf{v}_{i}^{(s)}))\;,$ (6)
$\mathbf{h}_{i}^{(s+1)}$ $=\operatorname{GRU}([F^{gcn}(\mathbf{v}_{i}^{(s)}),\mathbf{v}_{i}^{(s)}],\mathbf{h}_{i}^{(s)})\;,$ (7)
$\mathbf{v}_{i}^{(s+1)}$ $=\mathbf{v}_{i}^{(s)}+\text{MLP}(\mathbf{v}_{i}^{(s)},\mathbf{h}_{i}^{(s+1)})\;.$ (8)
In Eq. 5, the vertex features are extracted by trilinear interpolating $F^{0}_{\rm static}$ and $F^{t}_{\rm motion}$ at the vertex locations. They are then concatenated and passed through Eq. 6-8 to gradually update the vertex position, which aggregates the spatial information. Eq. 5-8 is repeated $K$ times, which we set to $2$ by default.
3.5 Two-Stage Weakly Supervised Training Using Only Keyframe Annotations
Creating high-quality annotations for an entire sequence is costly. As a result, many datasets provide annotations only for keyframes, such as end-diastolic (ED) and end-systolic (ES) phases. This rules out full supervision. To address this, we propose a two-stage weakly-supervised training scheme which makes training the reconstruction and motion branches feasible using only the available keyframe annotations.
3.5.1 Training the Reconstruction Branch
We start by training the reconstruction branch of Section 3.3. At each training iteration, we randomly select a labeled image $\mathbf{I}$ with ground truth segmentation and mesh annotation $L^{gt},\mathcal{M}^{gt}$. Let $L^{p},\mathcal{G}^{p}$ be our model’s segmentation and tetrahedra prediction. We minimize the loss
$\mathcal{L}_{shape}$ $(\mathbf{I})=\lambda_{cd}\mathcal{L}_{cd}(\textbf{MT}(\mathcal{G}^{p}),\mathcal{M}^{gt})$ (9)
$+\lambda_{sdf}\mathcal{L}_{sdf}(\mathcal{G}^{p},\mathcal{M}^{gt})+\lambda_{ce}\mathcal{L}_{ce}(L^{p},L^{gt})\;,$
where $\mathcal{L}_{cd}$ is the chamfer distance, $\mathcal{L}_{sdf}$ is an L1-loss used to supervise the predicted SDF values at tetrahedral grid vertices with the ground truth SDF value queried from the ground truth mesh. $\mathcal{L}_{ce}$ is the cross-entropy loss defined on the predicted segmentation map. MT denotes the marching tetrahedra algorithm [38] to convert $\mathcal{G}^{p}$ into surface mesh. $\lambda_{cd},\lambda_{sdf},\lambda_{ce}$ are set to $1.0,0.1,0.1$ respectively.
3.5.2 Training the Deformation Estimator
We then train the motion branch. At each iteration, we randomly select a sequence $\{\mathbf{I}^{t}\}_{t=0}^{T}$, and then choose an unlabeled frame $\mathbf{I}^{u}$ and a labeled frame $\mathbf{I}^{l}$ from it. We first predict $\mathcal{G}^{u}$ from $\mathbf{I}^{u}$. Next, assuming that $\mathbf{I}^{l}$ comprises a total of $D$ slices, we randomly pick 1 to $D$ 2D slices at random location to form $\hat{\mathbf{I}}^{l}$ to simulate the online few-slice case. $V_{\text{motion}}^{l}$ and $\hat{V}_{\text{motion}}^{l}$ are generated from $\mathbf{I}^{l}$ and $\hat{\mathbf{I}}^{l}$ using the AFA module. We write
$V_{\text{motion}}^{l}=AFA(F^{u}_{2d},F^{l}_{2d}),\hat{V}_{\text{motion}}^{l}=AFA(F^{u}_{2d},\hat{F}^{l}_{2d})\;.$ (10)
We use these 3D features to generate two different versions of deformed tetrahedra using Eq. 1: $\mathcal{G}^{u\rightarrow l}$ from full-slice encoded feature $V_{\text{motion}}^{l}$, and $\hat{\mathcal{G}}^{u\rightarrow l}$ from few-slice encoded feature $\hat{V}_{\text{motion}}^{l}$. The network is trained by minimizing
$\mathcal{L}_{motion}$ $(\mathbf{I}^{l},\mathbf{I}^{u})=\mathcal{L}_{distill}(V_{\text{motion}}^{l},\hat{V}_{\text{motion}}^{l})$ (11)
$+\mathcal{L}_{cd}(\text{MT}(\mathcal{G}^{u\rightarrow l}),\mathcal{M}^{l})+\mathcal{L}_{cd}(\text{MT}(\hat{\mathcal{G}}^{u\rightarrow l}),\mathcal{M}^{l}),$
[…]
where $\mathcal{M}^{l}$ is the ground truth mesh for $\mathbf{I}^{l}$ and sg denotes the stop-gradient operation. Minimizing the distillation loss $\mathcal{L}_{distill}$ ensures that the 3D information from full-slice $\mathbf{I}^{l}$ is successfully distilled into the few-slice encoded feature, while minimizing the chamfer losses enables the deformation model to infer accurate motion from the feature map encoded from any number of 2D slices, while retaining the ability to infer motion from complete 3D full-slice feature map. As a result, minimizing $\mathcal{L}_{motion}$ guarantees good reconstruction results for both full and few-slice inputs.
Once trained, our model can be used for new patients without retraining. Note that our training strategy actually enables the model to predict cardiac deformation starting from any phase, not just the end-diastolic (ED) phase. Additionally, by employing an aggressive sampling scheme, the model learns to reconstruct motion from a highly diverse set of input configurations. As a result, even though the online scenario is simulated during training, our model demonstrates strong generalization capability. It can robustly handle cases where $\mathbf{O}^{t}$ originates from a different sequence or where there is a scanning angle discrepancy relative to $\mathbf{I}^{0}$, consistently producing reliable motion predictions despite these challenges as will be shown in the experiment section.
4 Experiments
4.1 Experimental Settings
Table 1: Few-slice quantitative evaluation on unified dataset. We compared the predicted 3D meshes by deforming from ED frame to ES frame to the ground-truth 3D meshes. 1- / 5-Slice: Number of slices used to recover the motion. Due to space limit, we omit the standard deviation.
           1-Slice
Method GT-Free Optim-Free Repre. ACDC M&Ms M&Ms-2
CD (mm2) $\downarrow$ Dice (%) $\uparrow$ CD (mm2) $\downarrow$ Dice (%) $\uparrow$ CD (mm2) $\downarrow$ Dice (%) $\uparrow$
Myo LV RV Myo LV RV Myo LV RV Myo LV RV Myo LV RV Myo LV RV
4DMR [14] SDF 29.53 - - 75.63 - - 19.87 - - 76.12 - - 19.37 - - 74.11 - -
MR-Net [26] Mesh 28.61 40.70 46.15 77.43 82.12 73.33 14.59 16.46 25.11 77.59 80.35 73.94 15.61 16.68 25.23 76.29 82.66 76.23
Ours-Mesh [44] Mesh 17.42 25.11 42.33 80.14 84.90 74.32 11.26 14.17 23.19 79.88 82.11 75.96 11.98 15.11 22.63 77.66 84.97 77.99
Ours-SDF [46] SDF 18.65 26.45 43.78 79.55 83.43 74.00 12.16 14.88 23.91 79.09 81.62 74.23 12.15 15.72 23.34 77.01 84.37 77.15
Ours Tet 15.24 23.65 40.64 82.25 86.20 75.42 9.76 12.37 21.42 81.00 84.45 77.32 10.12 13.89 20.27 79.33 86.06 79.65
5-Slice
Method GT-Free Optim-Free Repre. ACDC M&Ms M&Ms-2
CD (mm2) $\downarrow$ Dice (%) $\uparrow$ CD (mm2) $\downarrow$ Dice (%) $\uparrow$ CD (mm2) $\downarrow$ Dice (%) $\uparrow$
Myo LV RV Myo LV RV Myo LV RV Myo LV RV Myo LV RV Myo LV RV
4DMR [14] SDF 24.56 - - 77.03 - - 16.16 - - 77.33 - - 15.52 - - 76.67 - -
MR-Net [26] Mesh 18.82 24.22 36.54 79.54 84.25 74.67 10.60 12.16 20.28 78.70 83.31 76.91 10.76 12.05 19.25 77.84 86.03 78.79
Ours-Mesh [44] Mesh 11.66 19.39 33.91 82.92 85.88 75.73 9.14 11.07 18.62 80.99 85.11 78.67 8.68 10.96 16.82 79.82 87.99 80.13
Ours-SDF [46] SDF 12.21 20.03 34.52 82.35 85.32 75.11 9.63 11.69 19.26 80.45 84.56 78.05 9.22 11.55 17.42 79.21 87.41 79.53
Ours Tet 10.22 18.31 32.61 84.38 87.32 77.23 7.93 9.72 17.39 82.83 86.76 79.27 7.30 9.96 15.58 81.24 89.37 81.48
Table 2: Clinical indexes evaluation on M&Ms dataset. 1- / 5- / Full-Slice denotes the number of slices used to recover motion.
  Method LVESV MAE $\downarrow$ LVEF MAE $\downarrow$ RVESV MAE $\downarrow$ RVEF MAE $\downarrow$
(ml) (%) (ml) (%)
1-Slice
MR-Net [26] 9.76 7.87 16.40 11.75
Ours-Mesh [44] 8.92 6.37 16.41 9.83
Ours-SDF [46] 9.35 6.82 17.04 10.21
Ours 7.51 4.83 15.03 8.51
5-Slice
MR-Net [26] 7.34 6.29 13.45 8.87
Ours-Mesh [44] 7.64 5.11 12.81 7.33
Ours-SDF [46] 7.93 5.32 13.12 7.59
Ours 6.86 4.30 11.88 6.55
Full-Slice
FFD [48] 23.53 14.95 24.06 15.87
dDemons [49] 33.97 22.11 63.03 38.52
MR-Net [26] 6.75 4.16 12.05 7.01
Ours-Mesh [44] 5.07 2.23 9.85 4.96
Ours-SDF [46] 5.22 2.34 9.98 5.01
Ours 4.84 2.04 9.65 4.77
4.1.1 Data
We train and validate our method on a unified dataset constructed from three large and diverse publicly available datasets: ACDC, M&Ms, and M&Ms-2 [19, 20, 21]. ACDC features CMR sequences of about 30 frames for 100 patients, each one covering at least one cardiac cycle and with a slice thickness of 5-10 mm. For these datasets, segmentation masks for the left ventricle (LV), right ventricle (RV), and myocardium (MYO) are provided, but only for the end-diastolic (ED) and end-systolic (ES) frames. M&Ms and M&Ms-2 are made of multi-center and multi-vendor collected CMR sequences for 375 and 360 subjects, respectively. While featuring similar resolutions and sequence lengths as ACDC dataset, these two datasets cover a broader range of cardiac pathologies. As in earlier studies [13, 14, 9, 3], we register and fit a Statistic Shape Model (SSM) [24, 5] to segmentation mask at the ED and ES frames using non-rigid image registration methods to generate cardiac meshes. The unified dataset comprises 835 CMR sequences, randomly split into 501 (60%), 167 (20%), and 167 (20%) sequences for training, validation, and testing, respectively.
We use two additional datasets for external evaluation to assess generalization in a clinical context where model retraining is not feasible. The first dataset was collected at the Center for Interventional MRI, University Hospital Lausanne (CHUV), designed with the target online scenario in mind and hereafter referred to as CHUV. For each subject, a stack of short-axis (SAX) breath-hold cine slices was acquired at rest on a 1.5T interventional system (Siemens Healthineers, Magnetom Aera) using an accelerated steady-state free precession (SSFP) sequence covering the entire left ventricle. In addition, single SAX cine slices were acquired at mid-ventricular levels with slight angulations relative to the full SAX stack. Because both heart shape and rate may vary during intervention, updating of the 3D heart model is often required to ensure precise intracardiac manipulations. To mimic heart-rate changes, five volunteers exercised inside the interventional MRI (bending and extending their legs in two parallel plastic gutters for 2–3 minutes), thereby increasing their heart rates by approximately 20–30 beats per minute. Once elevated, single SAX cine slices were reacquired at the same mid-ventricular levels as during rest. For real-time acquisition, the online sequences had lower spatial resolution than the offline short-axis CMR stacks. To evaluate full-sequence performance across the entire cardiac cycle beyond the keyframes, we further use the second external dataset, 4DM [14], which focuses on left ventricular myocardium reconstruction and provides mesh annotations for every frame. This dataset enables a rigorous evaluation of the proposed weakly supervised motion learning scheme.
4.1.2 Evaluation Metrics
For evaluation, we compare the predicted 3D mesh obtained by deforming the ED frame to the ES frame with the ground-truth 3D mesh at the ES frame. We use L2-norm Chamfer Distance (CD) to measure 3D reconstruction quality. Dice score is also reported to quantify the intra-slice segmentation results. Besides these quality measures, we also report the Mean-Absolute Error (MAE) of LV/RV ES Volume (LVESV/RVESV), LV/RV ejection fraction (LVEF/RVEF) derived from our predicted meshes. LVEF and RVEF are representative cardiac function indexes which could be used for clinical diagnosis.
4.1.3 Implementation Details
In all experiments, the slices are resampled by linear interpolation to a spacing of $1.25\times 1.25$ mm. From this, the AFA module constructs 3D volume with an effective spacing of $1.25\times 1.25\times 2$ mm. The AFA module uses a 8-head multi-head attention with learnable 3D position embedding. In the reconstruction branch, a modified nnU-Net [42] is used as the image encoder. It comprises five downsampling and upsampling blocks with channel sizes [32, 64, 128, 256, 320]. We use the same GCN structure as in [18] with $3$ layers and $128$ channels. The initial tetrahedra resolution is $128$. For the motion branch, we use the identical encoding network structure as the reconstruction branch. For the deformation model, the hidden dimension of the GCN and GRU are both set to $128$. We use SGD optimizer with an initial learning rate of 0.01, a weight decay of 3e53e-5, and momentum of $0.99$. We train all models for $300/150$ epochs for the reconstruction/motion learning stage. All experiments are conducted using one V100 GPU.
4.1.4 Baselines
We compare our method against several existing approaches for 4D shape reconstruction. For the online interventional setting, as noted earlier, most methods assume access to a complete CMR stack, making them unsuitable for the online few-slice setting. Therefore, we adapted several baselines so that they can operate under sparse inputs. 4DMR [14] is an implicit representation method that reconstructs myocardium mesh sequence by optimizing decoupled shape and motion latent codes at test time, minimizing the discrepancy between the predicted mesh and ground-truth contour points. We adapt it by restricting the optimization to incomplete contours extracted from the available few slices, while still using full contours for the initial frame. MR-Net [26] is a template-mesh-based approach that takes ground-truth contour points as input and leverages a hybrid PointNet + 3D CNN architecture to predict vertex deformations. Since MR-Net directly encodes point sets, it can naturally accommodate incomplete inputs, a scenario already considered in the original paper. Although originally designed for static reconstruction, we adapt it for motion estimation by setting the ED-frame mesh as the template.
Figure 2: 1-Slice Comparison. (Left) We predict 3D meshes by deforming from ED to ES frame. Color indicates the magnitude of point-to-surface error in $m$. (Right) Segmentation results on slices at the apex and base location. Note that 4DMR can only output motion for myocardium.
Figure 3: Motion sequence prediction results using different number of slices by deforming from ED frame to the target frame. Corresponding Volume-frame curve is given on the right. #1: NOR, #2: DCM.
In addition, we include two modified baselines derived from shape reconstruction methods: MeshDeformNet [44] and DeepCSR [46]. To adapt them for motion reconstruction, we integrated into both the same nnU-Net backbone, AFA module, deformation model, and training pipeline as our method. Since their main differences with our full framework lie in the shape representation, we denote them as Ours-Mesh and Ours-SDF.
For the traditional offline setting, in addition to the above methods, we further compare against B-spline Free Form Deformation (FFD) [48], diffeomorphic Demons (dDemons) [49], MR-Net w. nnU-Net, MulViMotion and DeepMesh. FFD and dDemons are classical registration algorithms that have been widely used in recent cardiac motion tracking studies [50, 51, 52]. MR-Net w. nnU-Net adopts the MR-Net architecture but replaces ground-truth contours with those predicted by nnU-Net, featuring a two-stage automated pipeline. MulViMotion and DeepMesh are mesh-based approaches that predict deformation fields via 3D CNNs, but only for the myocardium class. For fair comparison, we adopt their SAX-view-only versions as mentioned in the original papers.
Unless explicitly stated otherwise, all methods are jointly trained on the unified dataset.
4.2 Motion Estimation from a Few Slices
Table 3: Full-slice quantitative evaluation on unified dataset. We compared the predicted 3D meshes by deforming from ED frame to ES frame to the ground-truth 3D meshes.
  Method GT-Free Optim-Free Repre. ACDC M&Ms M&Ms-2
CD (mm2) $\downarrow$ Dice (%) $\uparrow$ CD (mm2) $\downarrow$ Dice (%) $\uparrow$ CD (mm2) $\downarrow$ Dice (%) $\uparrow$
Myo LV RV Myo LV RV Myo LV RV Myo LV RV Myo LV RV Myo LV RV
FFD [48] Mesh 21.84 38.85 51.04 69.46 79.56 72.04 42.01 63.23 88.77 62.84 76.65 66.23 40.79 66.42 95.66 63.13 79.80 76.02
dDemons [49] Mesh 15.72 22.97 35.85 77.06 86.87 73.60 32.16 44.39 61.15 65.83 79.87 65.09 34.08 48.77 66.22 67.25 83.44 72.39
4DMR [14] SDF 10.89 - - 80.03 - - 8.42 - - 82.33 - - 8.25 - - 82.67 - -
MR-Net [26] Mesh 14.02 19.21 24.07 80.02 86.35 78.93 8.86 10.03 16.86 79.40 86.90 79.73 8.39 7.24 12.99 80.72 87.86 80.60
MR-Net w. nnU-Net Mesh 15.11 20.45 26.13 79.66 85.38 76.45 9.88 10.25 18.60 78.56 85.00 77.21 9.83 9.24 13.35 79.56 85.89 78.31
MulViMotion [12] Mesh 17.21 - - 78.10 - - 13.75 - - 72.30 - - 12.33 - - 68.21 - -
DeepMesh [13] Mesh 16.42 - - 79.89 - - 9.24 - - 78.51 - - 9.20 - - 77.75 - -
Ours-Mesh [44] Mesh 6.80 11.84 23.41 86.62 88.11 84.74 4.26 5.08 13.92 84.09 87.12 83.73 4.15 5.12 11.04 85.01 88.67 86.65
Ours-SDF [46] SDF 7.21 12.65 22.78 86.95 88.43 84.00 4.44 5.26 14.21 83.87 86.91 83.77 4.30 5.44 11.32 84.55 88.37 86.10
Ours Tet 6.63 11.51 22.15 87.12 88.96 84.21 4.11 4.85 13.68 84.31 87.43 83.99 4.02 5.04 10.82 85.33 89.13 86.94
Table 4: Zero-shot evaluation results on public 4DM dataset. Check marks denote the dataset used for training. Oracle is trained and evaluated on 4DM.
  Method Training Dataset Myocardium CD (mm2) $\downarrow$
ACDC M&Ms M&Ms-2 4DM 1-Slice 5-Slice Full-Slice
FFD - - - - - - 11.51
dDemons - - - - - - 12.46
4DMR - 24.00 19.03 13.97
MR-Net - 15.87 11.56 7.97
DeepMesh - - - 8.32
Ours-Mesh - 13.31 10.32 5.97
Ours-SDF - 13.78 10.78 6.24
Ours - - - 25.43 19.42 13.81
- - - 16.91 12.54 8.23
- - - 17.88 13.03 8.83
- 11.33 9.08 5.51
Ours (Oracle) - - - 6.24 4.87 2.70
In this section we present results under the online few-slice scenario that has been the main motivation for this work. In Fig. 2, we show two representative motion reconstruction results using only the central slice as input: one from a normal (NOR) heart and another from a subject with dilated cardiomyopathy (DCM). Because of its stronger contractile function, the normal heart exhibits larger deformations than the DCM heart, making motion prediction more challenging, which is reflected by larger magnitudes in the error maps. Nevertheless, our method delivers accurate predictions in both cases. Furthermore, the segmentation results on slices at the apex and base location demonstrate that our model can provide reliable deformation estimation even in regions far from the slices used for motion inference. The complete reconstructed sequences together with the corresponding volume–time curves are shown in Fig. 3.
We report quantitative results on the whole test set in Tabs. 1 and at the top of Tab. 2. The first contains chamfer distances and dice scores and the second cardiac function index predictions. For consistency, we generated all these results using the central 1 or 5 slices even though slices at any position could be used, as shown in Section 4.4. TetHeart consistently outperforms our four baselines without requiring ground-truth segmentations or contours as input. Instead, it infers cardiac motion directly from raw CMR images without any time-consuming test-time optimization. As a result, it delivers fast inference—12 FPS on a single V100—with a potential for further acceleration by adopting a lightweight encoder architecture and leveraging model quantization techniques to meet the real-time constraints. Even when using a single slice, TetHeart delivers meaningful indices. All of this points toward its suitability for eventual clinical deployment.
Figure 4: Chamfer distance as a function of the number of slices used to capture the motion.
Although we incorporated into Ours-Mesh and Ours-SDF the same mechanisms for handling sparse features as in our complete method, the latter still delivers better results. This confirms that the tetrahedral representation we use is more effective at preserving spatial information within the object, making it a better choice for inferring 3D motion when only limited observations are available. To quantify this, we plot in Fig. 4 the performance of all three methods as a function of the number of slices used. The fewer, the greater the gap between our complete approach and the other two. Nevertheless, both variants still surpass prior approaches by a clear margin, demonstrating that design components such as the AFA module and our training strategy form a general and effective solution for handling incomplete slice scenarios.
The comparatively poor performance of 4DMR can be attributed to its reliance on a simple MLP architecture, which may lack the representational power to generalize well to the large and heterogeneous datasets considered in this study. Moreover, its test-time optimization strategy does not generalize well to severely under-sampled inputs. The weaker performance of MR-Net likely stems from the fact that it reconstructs meshes solely from contour points, thereby neglecting the rich information available in cine images.
4.3 Motion Estimation from Full Stacks
Figure 5: Full-Slice Comparison. (Left) We predict 3D meshes by deforming from ED to ES frame. Color indicates the magnitude of point-to-surface error in $m$. (Right) Segmentation results on slices at the apex and base location. Note 4DMR and DeepMesh can only output for myocardium.
Since we had to modify existing techniques to compare them to ours in the few-slice scenario, we now compare our method to the original baselines on the full set of slices they were designed to operate on and report the results in Tab. 3 and the bottom of Tab. 2. In Fig. 3 and 5, we show qualitative results on the same hearts as in Fig. 2, but now using all slices as input instead of only one.
Remarkably, even though the main driver of this work was operating on as few slices as possible, we also consistently outperform the baselines in the full-slice scenario. We attribute this to several factors. Consistent with the previous analysis, the relatively poor performance of MR-Net may be attributed to its neglect of the rich information available in cine images. Using predicted contours resulted in slightly inferior performance, indicating that such a two-stage automated pipeline is prone to error accumulation, a finding consistent with the conclusions in the original paper [26]. MulViMotion estimates voxel-wise motion fields and reconstructs meshes from warped segmentation maps, which is inherently less accurate than directly modeling meshes. Compared to DeepMesh, our model benefits from a shared encoding network, allowing the deformation branch to leverage informative features from the reconstruction branch. Moreover, our AFA module dynamically aggregates spatially adjacent features, further enhancing prediction quality.
4.4 External Evaluation: To Generalize without Retraining
We now test the generalizability of our model to the CHUV and 4DM datasets without retraining.
In Fig. 6, we show sequence predictions on CHUV with the spatial position of the real-time captured 2D slice visualized in the left. As can be seen, TetHeart reconstructs a plausible heart motion across the cardiac cycle under both rest and exercise situation. Moreover, the predicted volume–time curves further confirm the physiological plausibility of our motion reconstruction. In the exercise condition, the heart exhibits a shorter ES–ED recovery period and a reduced duration of each cardiac cycle, which aligns well with expectation.
As 4DM provides full-sequence annotations, but only for the myocardium. we evaluate the full-sequence prediction performance for the myocardium subclass, rather than focusing solely on the ES frame and report the results in Tab. 4. Here, check marks denote the dataset used for training. Oracle refers to training and testing directly on 4DM, serving as an upper bound for performance.
As shown in the table, our model demonstrates the best generalization abilities, achieving top performance in both few-slice and full-slice settings. In contrast, 4DMR exhibits poor generalization, likely due to significant distribution shifts across datasets and the limited representational capacity of its simple MLP architecture. While MulViMotion and DeepMesh show some generalization ability, their lower baseline performance causes them to lag behind our method.
Moreover, training with a larger and more diverse dataset improves generalization. And consistent with the findings from previous evaluations, representing the heart with tetrahedra enables more effective spatial information retention and exchange, particularly in scenarios with limited slice input.
Figure 6: Motion sequence prediction on the CHUV dataset. (Left) Spatial configuration of the initial SAX stack and the online observations. (Middle) Cardiac motion sequences for each subject under rest and exercise conditions. (Right) Volume–time curves.
4.5 Ablation Studies
Finally, we conduct ablation studies to evaluate our design choices. First, we examine the role of the AFA module. Simply stacking 2D features achieves comparable performance to AFA in the full-slice setting, but it fails to generalize to scenarios with variable numbers and positions of 2D slices. This highlights the necessity of a more flexible mechanism like AFA. Next, we evaluate the importance of spatial information in the AFA module. If spatial position encodings are removed from the attention computation, the model’s performance degrades moderately, indicating that spatial information is crucial for accurate feature aggregation. However, if we use only position encodings as the query and discard image features entirely, the performance drops substantially. This shows that both image features and spatial coordinates are essential and must work together for effective motion prediction. When training the motion branch, if we do not share parameters with the reconstruction branch and instead initialize the deformation branch randomly, the model performance drops significantly. This demonstrates the effectiveness of our parameter-sharing strategy: the 3D knowledge learned during reconstruction is successfully transferred to motion modeling.
Lastly, we ablate the impact of the distillation loss. Without it, the model achieves slightly better results in the full-slice setting, but the features encoded from sparse 2D inputs are less discriminative, leading to poorer performance under few-slice conditions. When the distillation loss is combined with our aggressive sampling strategy, the model achieves consistently stronger results across all settings, demonstrating the importance of this training scheme for improving robustness and generalization.
Table 5: Ablation study results on M&Ms dataset.
  Method M&Ms Myocardium CD (mm2) $\downarrow$
1-Slice 5-Slice Full-Slice
Ours 9.76 7.93 4.11
- AFA module - - 4.22
- position embedding 11.23 9.04 4.67
- image feature in query $Q$ 16.67 12.37 6.44
- shared encoding network 20.66 11.63 6.43
- distillation loss 10.34 8.12 4.03
5 Conclusion
We have proposed TetHeart, a unified framework for reconstructing heart motion from either full stacks of 2D MRI slices or from sparse sets, down to a single slice. We demonstrate, for the first time, that full 3D heart shape and motion can be inferred in sparse real-time CMR slices, making our approach suitable for guidance of interventional procedures. This breakthrough shifts cardiac reconstruction from retrospective analysis to live clinical use, laying the foundation for real-time digital twins of the beating heart. This capability not only has the potential to enable precise, image-guided interventions today, but also points to a future where digital twins of organs in motion guide therapies across medicine—adaptive, predictive, and personalized at the point of care.
[…]

### v2

@@
-End-to-End 4D Heart Mesh Recovery Across Full-Stack and Sparse Cardiac MRI
+End-to-End 4D Heart Recovery Across Full-Stack and Sparse Cardiac MRI
 Abstract
-Reconstructing cardiac motion from cine CMR sequences is critical for diagnosis, prediction, and intervention.
-Existing methods rely on complete CMR stacks to infer full heart motion, limiting their utility in intra-procedural scenarios where only sparse observations are available.
-We present TetHeart, the first end-to-end framework that unifies full 4D multi-structure heart mesh recovery from both offline full-stack acquisitions and intra-procedural sparse-slice observations.
-Our method leverages deep deformable tetrahedra, an explicit–implicit hybrid representation, to capture shape and motion in a coherent space shared across cardiac structures.
-It is initialized from high-quality pre-procedural or offline-acquired full stacks to build detailed, patient-specific heart meshes, which can then be updated using whatever slices are available, from full stacks down to a single slice.
-We further incorporate several key innovations: (i) an attentive mechanism for slice-adaptive 2D–3D feature assembly that dynamically integrates information from arbitrary numbers of slices at any position, combined with a distillation strategy from full-slice to sparse-slice settings to ensure accurate reconstruction under extreme sparsity; and (ii) a two-stage weakly supervised motion learning scheme requiring only keyframe (e.g., ED and ES) annotations.
-Trained and validated on three large public datasets and externally evaluated zero-shot on additional private interventional and public CMR datasets, TetHeart achieves state-of-the-art accuracy and strong generalization in both pre- and intra-procedural settings.
-{IEEEkeywords}
-cardiac magnetic resonance (CMR), deformable tetrahedra, explicit–implicit hybrid representation, shape and motion modeling, sparse observation.
+Reconstructing cardiac motion from CMR sequences is critical for diagnosis, prognosis, and intervention.
+Existing methods rely on complete CMR stacks to infer full heart motion, limiting their applicability during intervention when only sparse observations are available.
+We present TetHeart, the first end-to-end framework for unified 4D heart mesh recovery from both offline full-stack and intra-procedural sparse-slice observations.
+Our method leverages deformable tetrahedra to capture shape and motion in a coherent space shared across cardiac structures.
+Before a procedure, it initializes detailed, patient-specific heart meshes from high-quality full stacks, which can then be updated using whatever slices can be obtained in real-time, down to a single one during the procedure.
+TetHeart incorporates several key innovations: (i) an attentive slice-adaptive 2D–3D feature assembly mechanism that integrates information from arbitrary numbers of slices at any position; (ii) a distillation strategy to ensure accurate reconstruction under extreme sparsity; and (iii) a weakly supervised motion learning scheme requiring annotations only at keyframes, such as the end-diastolic and end-systolic phases.
+Trained and validated on three large public datasets and evaluated zero-shot on additional private interventional and public datasets without retraining, TetHeart achieves state-of-the-art accuracy and strong generalization in both pre- and intra-procedural settings.
 1 Introduction
-The heart is one of the most vital organs in the human body [1], and accurate modeling of its shape and motion is central to numerous medical applications ranging from diagnosis to therapy planning [2, 3, 4].
-Cardiac magnetic resonance (CMR) offers rich temporal imaging of cardiac anatomy and has become an indispensable modality for both shape [5, 6, 7, 8, 9] and motion [10, 11, 12, 13, 14] modeling.
-However, to this day, such modeling remains confined to offline analysis and pre-operative planning.
-Existing techniques [5, 6, 7, 8, 9, 10, 11, 12, 13, 14] cannot be applied during an intervention, as they require full-stack CMR data that cannot be acquired in real time.
-In the intervention room, only a handful of real-time slices can be captured, usually at lower resolution [15, 16, 17], a limitation also confirmed in routine practice at our collaborating university hospital.
-This scenario motivates us to move beyond the current state of the art and recover full 3D shape and motion from sparse and partial observations.
+Modeling cardiac shape and motion is a vital 4D reconstruction problem with direct implications for medical imaging, motion analysis, and image-guided intervention [39, 22, 33, 32].
+Cardiac magnetic resonance (CMR) provides high-quality temporal anatomy and has been widely used for both shape reconstruction [2, 51, 14, 50, 48] and motion estimation [34, 26, 27, 28, 52].
+However, as shown in Fig.
+1 (left), existing methods rely on full volumetric CMR stacks and are thus restricted to offline analysis and pre-operative planning.
+In interventional settings, as shown in Fig.
+1 (right), only a handful of low-resolution slices can be acquired in real time [16, 29, 36], a limitation we observed in practice at our collaborating university hospital.
+This makes current 4D modeling pipelines unusable during intervetions, which motivated us to go beyond the current state of the art and to recover full 3D shape and motion from sparse and partial observations.
 This will enable real-time tracking of patient-specific heart dynamics, with a view to delivering guidance to interventional teams.
-To this end, we introduce TetHeart, a unified framework for reconstructing the full 4D motion of multiple cardiac structures from either complete CMR stacks or intra-procedural sparse data.
-As in Fig.
-1, TetHeart first leverages high-quality pre-operative or offline-acquired CMR stacks to build a detailed, patient-specific heart mesh, and then updates it in an end-to-end manner using any available slices, from a complete stack down to a single slice.
-To make this possible, TetHeart builds on deep deformable tetrahedra [18], an explicit–implicit hybrid representation that unites the optimization efficiency of signed distance functions with the geometric manipulability of meshes, forming the foundation for jointly modeling shape and motion in a coherent space shared across structures.
-On top of this, our first core contribution is an attentive mechanism for slice-adaptive 2D-3D feature assembly, which enables the model to handle sparse input by dynamically integrating information from arbitrary slices at arbitrary positions.
-Additionally, we introduce a distillation strategy from full-slice to sparse-slice settings, ensuring accurate reconstruction under extreme input sparsity.
-Our second contribution is a two-stage weakly supervised motion learning scheme that requires only keyframe (e.g., ED and ES) annotations, making it feasible to train TetHeart across diverse clinical environments.
-The core idea is to first train a static shape reconstruction branch on labeled frames, and then repurpose it to initialize shapes for unlabeled frames, from which motion dynamics are recovered on the labeled frames to provide supervision.
+Figure 1: Difference between the standard offline scenario and the intervention room setting for cardiac motion reconstruction.
+To this end, we introduce TetHeart, a unified framework for 4D cardiac motion reconstruction that operates consistently across both full-stack and sparse-slice CMR inputs, the latter being what can be acquired in real-time during an intervention.
+As shown in Fig.
+2, TetHeart first constructs a patient-specific mesh representation from pre-operative data and then dynamically updates it during procedures using only the slices that can be acquired in real-time, down to a single one.
+TetHeart builds on deep marching tetrahedra [40], which combines the optimization efficiency of signed distance fields with the geometric flexibility of tetrahedral meshes, providing a coherent space for joint shape–motion learning.
+Given this, our main contributions are as follows:
+  • We propose a slice-adaptive 2D–3D feature fusion mechanism that allows our networks to dynamically aggregate features from arbitrarily located slices, enabling robust reconstruction across varying input sparsity.
+  • We introduce a full-to-sparse distillation strategy that transfers knowledge from dense to sparse regimes, ensuring accurate reconstruction under extreme input sparsity.
+  • We develop a weakly supervised motion learning scheme that requires annotations only at key cardiac phases (end-diastolic and end-systolic) and learns motion through temporal consistency.
+This is essential in practice because it makes the training feasible and affordable across diverse clinical environments.
+  • We demonstrate that a single model trained once generalizes seamlessly across offline and online scenarios, achieving state-of-the-art results on ACDC [5], M&Ms [9], and M&Ms-2 [25].
+It also shows strong generalizability by further validating on a privately collected interventional CMR dataset and the public 4DM [52] dataset without retraining, demonstrating its potential for deployment in real-world clinical environments.
+In short, the combination of our feature fusion mechanism and training scheme yield a novel unified framework that enables consistent reconstruction and motion inference across a wide spectrum of observational sparsity.
+As a result, TetHeart can operate both offline and online, thus enabling comprehensive, patient-specific motion tracking across the entire clinical workflow.
+We achieve state-of-the-art accuracy for offline use with full-stack CMR, and perhaps more importantly, we unlock online use in intra-procedural scenarios where only sparse observations are available, which cannot be handled using current methods.
+We will make our code publicly available.
 Figure 2: TetHeart reconstructs cardiac motion from observed 2D slices.
@@
 For each subsequent time step $t$, the observed 2D slices $\mathbf{O}^{t}$ are encoded via the AFA module with cross-attention to obtain a volumetric motion feature $V_{\text{motion}}^{t}$, which is then used to recover motion by predicting offsets that deform $\mathcal{G}^{0}$.
-Weights are shared between the static reconstruction branch and dynamic deformation branch.
-By representing the heart in a shared deformable space, TetHeart preserves spatial alignment between structures while adapting to varying sparsity levels.
-This offline/online capability enables comprehensive, patient-specific motion tracking across the entire clinical workflow: not only achieving state-of-the-art accuracy for offline use with full-stack CMR, but also, more importantly, unlocking online use in intra-procedural scenarios where only sparse observations are available.
-We conducted extensive experiments to train and validate our model on three large public datasets: ACDC [19], M&Ms [20], and M&Ms-2 [21].
-Remarkably, Remarkably, with a single set of weights learned from only end-diastolic (ED) and end-systolic (ES) annotations, our method generalizes seamlessly across both online few-slice and offline full-slice settings, consistently outperforming existing approaches not only on geometric metrics (Chamfer Distance and Dice) but also on clinically relevant indexes related to ejection fraction.
-We further performed external evaluations without retraining to assess generalizability.
-First, we tested the method on a privately collected dataset tailored for the online interventional MRI scenario at our collaborating cardiac CMR center, demonstrating its potential for deployment in real-world clinical environments.
-In addition, we evaluated the model on the public 4DM dataset [14], which provides complete temporal annotations of the myocardium, thereby demonstrating that the proposed weakly supervised motion learning scheme enables accurate reconstruction across the entire cardiac cycle.
-Together, these results highlight the effectiveness, robustness, and clinical promise of TetHeart.
-We will make our code public.
+Weights are shared between the static and dynamic branch.
 2 Related Work
+We first discuss current approaches to recovering static and dynamic hearts, and then the imaging modality we focus on.
 2.1 Static Cardiac Shape Modeling
 Reconstructing the heart from 3D medical images holds significant clinical value.
-However, cardiac MR (CMR) images often exhibit low through-plane resolution, which makes accurate reconstruction a challenge.
-Over the years, many approaches to addressing it have been proposed [22, 5, 23, 24, 25, 26, 27, 28, 29, 9] .
-They can be categorized into traditional optimization-based techniques and deep learning-based methods.
-2.1.1 Optimization-Based Approaches
-They usually rely on a two-stage pipeline.
-Contours are first extracted from the CMR images—either manually or using a segmentation algorithm—and then a template mesh is optimized to fit these contours.
-For example, the approach of [23] starts from a predefined tubular initial mesh that is deformed to minimize a contour-matching loss.
-More recently, a 3D active shape model (3D-ASM) and an image intensity model was used in [29] to align initial meshes with ground-truth contours.
-While these approaches can produce good results, they also have drawbacks.
-As they rely on test-time optimization, the reconstruction process is often time-consuming, taking from tens of seconds to several minutes per mesh.
-When applied to a full image sequence, this may extend to over an hour, making them impractical for real-time applications.
-Moreover, these methods’ performance has only been evaluated using ground-truth contours, which require labor-intensive manual annotation.
-What their performance is when using predicted contours instead remains an open question.
-2.1.2 Deep Learning-Based Approaches
-Deep learning-based methods leverage prior knowledge learned from large datasets to accelerate mesh reconstruction.
-For instance, in [28], sets of contours are treated as sparse point cloud and decoded into 3D meshes using point- and graph-convolutions.
-Similarly, in [26], contour points are embedded into a 3D volume and a hybrid architecture made of 3D CNNs and GCNs is used to progressively deform a template mesh.
-These methods increase speed but also suffer from several limitations.
-Because meshes are reconstructed solely from extracted contours or points, the rich appearance information contained in cine images is not fully exploited.
-Furthermore, as contour extraction is not necessarily perfect, such two-states approaches can result in error accumulation [26].
+However, CMR images often exhibit low through-plane resolution, making accurate reconstruction challenging.
+Current solutions [42, 2, 45, 15, 1, 10, 47, 4, 18, 48] use either traditional optimization-based techniques or deep learning-based methods.
+Optimization-Based Approaches usually rely on a two-stage pipeline, often starting from contours manually extracted from the CMR images.
+For example, [45] starts from a predefined tubular mesh that is deformed to minimize a contour-matching loss and, in [18], a 3D active shape model and an intensity model are used to align initial meshes with GT contours.
+Even though this works, relying on test-time optimization makes the reconstruction time-consuming, taking from tens of seconds to several minutes per frame.
+When applied to a full image sequence, this may extend to over an hour, making it impractical for real-time applications.
+Moreover, it has only been evaluated using GT contours, requiring labor-intensive manual annotation.
+Deep Learning-Based Approaches leverage prior knowledge learned from large datasets to accelerate mesh reconstruction.
+For instance, in [4], contours are treated as sparse point cloud and decoded into 3D meshes using point- and graph-convolutions.
+Similarly, in [10], contours are embedded into a 3D volume and a model made of 3D CNNs and GCNs is used to progressively deform a template mesh.
+These methods are much faster but also have drawbacks.
+Reconstructing meshes solely from extracted contours or points fails to exploit the rich appearance information contained in cine images [48].
+Furthermore, as contour extraction is not necessarily perfect, such two-step approaches are subject to error accumulation [10].
 2.2 Dynamic Cardiac Motion Modeling
@@
 However, this would fail to exploit temporal consistency and make it difficult to establish inter-frame correspondences.
-Thus a number of motion-based approaches that compute the deformation from frame to frame have been proposed.
-2.2.1 Motion from Full-Slice Offline MRI Sequences
+Thus, several motion-based approaches that compute the deformation from frame to frame have been proposed.
 When full CMR stacks are available, recent deep learning-based approaches [26, 27, 28, 52, 51] can be used.
 MulViMotion [27] trains 3D CNNs to estimate left ventricular myocardial motion by predicting voxel-based deformation fields from CMR data, and then building 3D meshes by warping segmentation masks.
-DeepMesh [13] decouples mesh reconstruction from motion estimation, and uses 3D CNNs and bilinear interpolation to predict per-vertex deformation.
+DeepMesh [28] decouples mesh reconstruction from motion estimation.
+It uses 3D CNNs and interpolation to predict per-vertex deformation.
 4DMR [52] models the myocardium using implicit representation.
-Decoupled shape and motion latent codes are optimized during test time to predict an initial shape and the backward motion fields.
-We use MulViMotion, DeepMesh and 4DMR as baselines because they are among the best currently available.
+Decoupled shape and motion latent codes are optimized at test time to predict a source shape and motion fields.
+We use MulViMotion, DeepMesh and 4DMR as baselines because they are among the best current methods.
 Unfortunately, these approaches also suffer from several limitations.
-First, some of them focus only on myocardium, neglecting other cardiac structures, preventing comprehensive heart assessment.
+First, some of them focus solely on the myocardium, neglecting other cardiac structures and preventing comprehensive heart assessment.
 Second, optimization-based methods like 4DMR are time-consuming at inference time.
-And, finally and most importantly, they require full scans that cannot be acquired at sufficient frame rate, which precludes intra-procedural use as discussed below.
-2.2.2 Motion from Sparse Online MRI Sequences
-CMR imaging has a reputation for being a ”slow” modality.
+Finally and most damagingly, they require full scans that cannot be acquired at sufficient frame rate, which precludes intra-procedural use as discussed below.
+2.3 Real-Time MR Imaging
+MRI has a reputation for being slow.
 Acquiring a full MRI stack involves sequential slice acquisition over a 4–6 heartbeat breath-hold, which can take several minutes [29, 36].
-While this is feasible for pre-interventional planning, it is not suitable for intra-procedural settings.
-Recently, real-time MRI (RT-MRI) has gained favor due to its ability to capture dynamic processes without the need for gating, synchronization, or repetition [16].
-Its applications include complex electrophysiology procedures [30], congenital heart disease [31], and other cardiac interventions [15, 32].
+While this is feasible for pre-interventional planning, but impractical for intra-procedural work.
+Recently, real-time MRI (RT-MRI) has gained favor due to its ability to capture dynamic processes without gating, synchronization, or repetition [29].
+Its applications include complex electrophysiology procedures [30], congenital heart disease [17], and cardiac interventions [16, 49].
 RT-MRI provides excellent soft-tissue contrast for assessing interventions like ablations [6, 36] and allows for continuous visualization and tracking of devices, and it is free of ionizing radiation [30, 8].
-Nevertheless, although RT-MRI enables rapid and continuous image acquisition, it still involves a fundamental tradeoff between spatial resolution, temporal resolution, artifacts and reconstruction latency [16].
-Therefore, achieving a sufficient temporal resolution (e.g., 20–25 Hz) to be used during invasive procedures, can only be done at the cost of acquiring a limited number of 2D slices at lower spatial resolution [15, 32, 34, 16, 17].
+Although RT-MRI enables rapid continuous image acquisition, it still involves a fundamental tradeoff between spatial resolution, temporal resolution, artifacts and reconstruction latency [29].
+Therefore, achieving a sufficient temporal resolution (e.g., 20–25 Hz) for use during invasive procedures, can only be done at the cost of acquiring a limited number of 2D slices at lower spatial resolution [16, 49, 8, 29, 36].
 We are not aware of any current method that can reconstruct 3D cardiac motion using this kind of data.
@@
 3 Method
-In this section, we introduce TetHeart, our framework for accurate 4D mesh reconstruction from either full CMR scans or sparse sets of 2D slices, as depicted by Fig 1.
-We first introduce our heart presentation that is based on deep deformable tetrahedra [18], an explicit–implicit hybrid representation that unites the optimization efficiency of signed distance functions with the geometric manipulability of meshes.
-We then introduce our Attentive 2D–3D Feature Assembler (AFA) module, which enables effective feature encoding from 2D slices with varying numbers and locations.
-Finally, we present our weakly supervised training scheme that makes it possible to train given only a limited number of annotations.
+We now present TetHeart, a framework for accurate 4D cardiac shape and motion reconstruction from either full CMR volumes or sparse 2D slice sets, as shown in Fig.
+2.
+We first describe our heart representation based on deep marching tetrahedra [40], an explicit–implicit hybrid model combining the optimization efficiency of signed distance fields with the geometric flexibility of tetrahedral meshes.
+Next, we introduce the Attentive 2D–3D Feature Assembler (AFA) we designed to adaptively aggregate features from arbitrary numbers of 2D slices at varying spatial locations.
+Finally, we detail our weakly supervised training scheme, enabling effective motion learning from limited annotations available only at key cardiac phases.
 3.1 Formalization
-To model deforming hearts, we use the deep deformable tetrahedra [18] formalism.
-We discretize the 3D space using a deformable tetrahedral grid, in which each vertex possesses an SDF value, denoted as $\mathcal{G}=(\{\mathbf{v}_{i}\},\{s_{i}\},\mathcal{T})$, where $\mathbf{v}_{i}$ and $s_{i}$ are the vertices and their corresponding SDF values in $\mathcal{T}$, the set of all tetrahedrons.
+To model deforming hearts, we use the deep marching tetrahedra formalism.
+We discretize the 3D space using a deformable tetrahedral grid, in which each vertex possesses an SDF, denoted as $\mathcal{G}=(\{\mathbf{v}_{i}\},\{s_{i}\},\mathcal{T})$, where $\mathbf{v}_{i}$ and $s_{i}$ are the vertices and their corresponding SDFs in $\mathcal{T}$, the set of all tetrahedrons.
 The explicit surface can then be extracted using the differentiable marching tetrahedra algorithm [13].
 Given an initial 3D heart model $\mathcal{G}^{0}$ reconstructed from the end-diastolic (ED) frame $\mathbf{I}^{0}\in\mathbb{R}^{D\times H\times W}$, our goal is to estimate a deformed model $\mathcal{G}^{t}$ at each time instant $t$, based on subsequent observations $\mathbf{O}^{t}$ that reflect the ongoing cardiac motion.
-We do this by updating the vertices of $\mathcal{G}^{0}$.
-We write
+We do this by updating the vertices of $\mathcal{G}^{0}$:
 $\mathcal{G}^{t}=(\{\mathbf{v}_{i}^{t}\},\{s_{i}\},\mathcal{T})\;,\mathbf{v}_{i}^{t}=\mathcal{D}(\mathbf{v}_{i}^{0},\mathbf{O}^{t})\;,$ (1)
@@
     Online.
-For intra-procedural use, $\mathbf{I}^{0}\in\mathbb{R}^{D\times H\times W}$ is a full-slice image acquired before the intervention, taking as much time as necessary.
+For intra-procedural use, $\mathbf{I}^{0}\in\mathbb{R}^{D\times H\times W}$ is a full-slice image acquired before the intervention, taking as much time as needed.
 However, it has to be possible to acquire $\mathbf{O}^{t}$ at sufficient frame rate and, consequently, it can only comprise a small set—from 1 to 3—of 2D slices at lower spatial resolution and at arbitrary locations [16].
@@
 3.2 Attentive 2D-3D Feature Assembler
-Because the observations can be either a full CMR stack, a sparse set of slices, or anything in between, we need a mechanism to extract useful features in all cases.
-To this end, we developed the Attentive 2D-3D Feature Assembler (AFA) module that relies on an attention mechanism that takes $\mathbf{I}^{0}$ and $\mathbf{O}^{t}$ as input and produces features.
-Let us write $\mathbf{O}^{t}$ as $\{\mathbf{I}^{t}_{s}\in\mathbb{R}^{H\times W}|s=1,\ldots,S\}$, a collection of 2D slices where $S$ is the total number of slices.
-With this notation, $\mathbf{I}^{0}$ can be written as $\{\mathbf{I}^{0}_{d}\in\mathbb{R}^{H\times W}|d=1,\ldots,D\}$.
+Because the observations can be either a full stack, a sparse set of slices, or anything in between, we need a mechanism to extract useful features in all cases.
+To this end, we developed the Attentive 2D-3D Feature Assembler (AFA) module that takes $\mathbf{I}^{0}$ and $\mathbf{O}^{t}$ as input and produces features.
+Let us write $\mathbf{O}^{t}$ as $\{\mathbf{I}^{t}_{s}\in\mathbb{R}^{H\times W}|s=1,\ldots,S\}$, a collection of 2D slices where $S$ is the number of slices.
+With this, $\mathbf{I}^{0}$ can be written as $\{\mathbf{I}^{0}_{d}\in\mathbb{R}^{H\times W}|d=1,\ldots,D\}$.
 A full-slice observation is simply one where $S=D$.
@@
 Motion Features.
-A naive way to use an attention mechanism to construct motion features would be to treat the $F^{0}_{2d}$ as the query and the $F^{t}_{2d}$ as both keys and values.
+A naive way to use attention to construct motion features would be to treat the $F^{0}_{2d}$ as the query and the $F^{t}_{2d}$ as both keys and values.
 However, this would be inefficient because global attention aims at capturing long-range dependencies, which is not needed to model the inherently local nature of anatomical motion.
 Also, this requires computing full pairwise relationships between all query and key positions, which is computationally expensive when using many slices and high-resolution feature maps.
-To avoid this, we take our inspiration from the Swin Transformer [39] and use instead a variant of attention that preserves locality and significantly reducing computational complexity.
+To avoid this, we take our inspiration from the Swin Transformer [24] and use instead a variant of attention that preserves locality and significantly reduces computational complexity.
 For each voxel in $F^{0}_{2d}$, we first identify $k_{1}$ spatially nearest 2D slices, or all slices if there are only $k_{1}$ or less.
@@
 This can be written as
-$V^{t}_{\rm motion}=AFA(F^{0}_{2d},F^{t}_{2d})=\text{MHAttention}(Q,K,V)\;,$ (2)
+$V^{t}_{\rm motion}=AFA(F^{0}_{2d},F^{t}_{2d})=\text{MHAttention}(Q,K,V)\;,$
 $\text{where }Q=F^{0}_{2d},~K=V=\text{NN-Select}_{F^{0}_{2d},k_{1},k_{2}}(F^{t}_{2d})\;,$
@@
 MHAttention denotes a Multi-Head Attention [43] operation.
-A learnable position embedding is used to encode the spatial information into the calculation of attention.
+A learnable position embedding is used to encode the spatial information.
 In practice, we set $k_{1}=3$ and $k_{2}=9$.
@@
 While our AFA module effectively encodes features from varying 2D slice configurations, a persistent challenge remains in the online scenario, where the limited number of slices can hinder accurate 3D motion reconstruction.
-To alleviate this issue, we introduce a distillation loss that encourages the model to efficiently transfer knowledge from full-slice inputs to partial-slice inputs during training.
+To alleviate this issue, we introduce a distillation loss that encourages the model to transfer knowledge from full-slice inputs to partial-slice inputs during training.
 In short, we predict two versions of motion features from full-slice observation and its randomly sampled subset.
@@
 This can be similarly written as
-$V^{0}_{\rm static}=AFA(F^{0}_{2d},F^{0}_{2d})=\text{MHAttention}(Q,K,V)\;,$ (3)
+$V^{0}_{\rm static}=AFA(F^{0}_{2d},F^{0}_{2d})=\text{MHAttention}(Q,K,V)\;,$
 $\text{where }Q=F^{0}_{2d},~K=V=\text{NN-Select}_{F^{0}_{2d},k_{1},k_{2}}(F^{0}_{2d})\;.$
@@
 Features from the last level of the upsample stream are taken as the final features, which we denoted as $F^{0}_{\rm static}$ and $F^{t}_{\rm motion}$, respectively.
-Note that the initial 2D encoding block and the subsequent 3D convolutions are shared for both static and motion feature extraction.
-This parameter sharing strategy enabled by our unified design allows motion prediction to benefit from the rich 3D spatial information learned in the static reconstruction task.
-Consequently, it facilitates more accurate motion estimation—an advantage overlooked by some previous works [13, 14].
-As demonstrated in the experiments, this architecture leads to superior motion reconstruction performance.
+Note that the initial 2D encoding block and subsequent 3D convolutions are shared for static and motion feature extraction.
+This unified design allows motion prediction to benefit from the rich 3D spatial information learned in the static reconstruction task, facilitating more accurate motion estimation—an advantage overlooked by some previous works [28, 52].
+As shown in the experiments, this design leads to superior motion reconstruction performance.
 3.3 Initial Heart Model Reconstruction
-Given the $F^{0}_{\rm static}$ features computed as discussed above from a full-slice CMR image $\mathbf{I}^{0}$, our first task is to reconstruct the initial tetrahedral mesh $\mathcal{G}^{0}$.
-To this end, we start from an initial tetrahedral grid $\mathcal{G}^{init}$ obtained by uniformly sampling a unit cube.
+Given $F^{0}_{\rm static}$ computed as discussed above, our first task is to reconstruct the initial tetrahedral mesh $\mathcal{G}^{0}$.
+To this end, we start from tetrahedral grid $\mathcal{G}^{init}$ obtained by uniformly sampling a unit cube.
 We then trilinearly interpolate $F^{0}_{\rm static}$ at the vertices of $\mathcal{G}^{init}$.
-An SDF value and an offset for each vertex is then predicted using a GCN to create $\mathcal{G}^{0}$.
+An SDF value and an offset for each vertex is predicted using a GCN to create $\mathcal{G}^{0}$.
 We write
@@
 $F^{0}_{\rm static}(\mathbf{v})$ denotes feature extraction at $\mathbf{v}$ from $F^{0}_{\rm static}$ with trilinear interpolation, and $\cdot$$\cdot$] denotes concatenation.
-To enhance the network’s spatial sensitivity, the vertex location is appended to the feature vector before it is passed to the GCN.
-For datasets with $C$ classes, we simply set the output channel of the GCN layer to $C\times 4$, 3 for deformation, and 1 for SDF value.
-This yields a separate tetrahedral mesh for each class.
+To enhance spatial sensitivity, the vertex location is appended before it is passed to the GCN.
+For datasets with $C$ classes, we set the output channel to $C\times 4$, 3 for deformation and 1 for SDF.
 Following prior works [46, 21, 7, 12], our network also has an auxiliary segmentation prediction branch during training time.
@@
 In Eq.
-5, the vertex features are extracted by trilinear interpolating $F^{0}_{\rm static}$ and $F^{t}_{\rm motion}$ at the vertex locations.
+3, the vertex features are extracted by trilinear interpolating $F^{0}_{\rm static}$ and $F^{t}_{\rm motion}$ at their locations.
 They are then concatenated and passed through Eq.
-6-8 to gradually update the vertex position, which aggregates the spatial information.
-Eq.
-5-8 is repeated $K$ times, which we set to $2$ by default.
+4-6 to gradually update the vertex position.
+Eqs.
+3-6 are repeated twice.
 3.5 Two-Stage Weakly Supervised Training Using Only Keyframe Annotations
 Creating high-quality annotations for an entire sequence is costly.
-As a result, many datasets provide annotations only for keyframes, such as end-diastolic (ED) and end-systolic (ES) phases.
+Therefore, many datasets provide annotations only for keyframes, such as end-diastolic (ED) and end-systolic (ES) phases.
 This rules out full supervision.
-To address this, we propose a two-stage weakly-supervised training scheme which makes training the reconstruction and motion branches feasible using only the available keyframe annotations.
-3.5.1 Training the Reconstruction Branch
+To address this, we propose a two-stage weakly-supervised pipeline, which makes training the reconstruction and motion branches feasible using only the available keyframe annotations.
+Training the Reconstruction Branch.
 We start by training the reconstruction branch of Section 3.3.
@@
 $\lambda_{cd},\lambda_{sdf},\lambda_{ce}$ are set to $1.0,0.1,0.1$ respectively.
-3.5.2 Training the Deformation Estimator
+Training the Deformation Estimator.
 We then train the motion branch.
-At each iteration, we randomly select a sequence $\{\mathbf{I}^{t}\}_{t=0}^{T}$, and then choose an unlabeled frame $\mathbf{I}^{u}$ and a labeled frame $\mathbf{I}^{l}$ from it.
+At each iteration, we randomly select a sequence $\{\mathbf{I}^{t}\}_{t=0}^{T}$, and choose an unlabeled frame $\mathbf{I}^{u}$ and a labeled frame $\mathbf{I}^{l}$ from it.
 We first predict $\mathcal{G}^{u}$ from $\mathbf{I}^{u}$.
@@
 We write
-$V_{\text{motion}}^{l}=AFA(F^{u}_{2d},F^{l}_{2d}),\hat{V}_{\text{motion}}^{l}=AFA(F^{u}_{2d},\hat{F}^{l}_{2d})\;.$ (10)
+$V_{\text{motion}}^{l}=AFA(F^{u}_{2d},F^{l}_{2d}),\hat{V}_{\text{motion}}^{l}=AFA(F^{u}_{2d},\hat{F}^{l}_{2d})\;.$
 We use these 3D features to generate two different versions of deformed tetrahedra using Eq.
@@
 where $\mathcal{M}^{l}$ is the ground truth mesh for $\mathbf{I}^{l}$ and sg denotes the stop-gradient operation.
-Minimizing the distillation loss $\mathcal{L}_{distill}$ ensures that the 3D information from full-slice $\mathbf{I}^{l}$ is successfully distilled into the few-slice encoded feature, while minimizing the chamfer losses enables the deformation model to infer accurate motion from the feature map encoded from any number of 2D slices, while retaining the ability to infer motion from complete 3D full-slice feature map.
+Minimizing the distillation loss $\mathcal{L}_{distill}$ encourages features from few-slice input $\hat{\mathbf{I}}^{l}$ to align with the one from full-slice input $\mathbf{I}^{l}$, while minimizing the chamfer losses enables the deformation model to infer accurate motion from the feature map encoded from any number of 2D slices, while retaining the ability to infer motion from complete 3D full-slice feature map.
 As a result, minimizing $\mathcal{L}_{motion}$ guarantees good reconstruction results for both full and few-slice inputs.
 Once trained, our model can be used for new patients without retraining.
-Note that our training strategy actually enables the model to predict cardiac deformation starting from any phase, not just the end-diastolic (ED) phase.
-Additionally, by employing an aggressive sampling scheme, the model learns to reconstruct motion from a highly diverse set of input configurations.
+By employing an aggressive sampling scheme, the model learns to reconstruct motion from a highly diverse set of input configurations.
 As a result, even though the online scenario is simulated during training, our model demonstrates strong generalization capability.
-It can robustly handle cases where $\mathbf{O}^{t}$ originates from a different sequence or where there is a scanning angle discrepancy relative to $\mathbf{I}^{0}$, consistently producing reliable motion predictions despite these challenges as will be shown in the experiment section.
+It can robustly handle cases where $\mathbf{O}^{t}$ originates from a different sequence or where there is a scanning angle discrepancy relative to $\mathbf{I}^{0}$, consistently producing reliable motion predictions as will be shown in the experiment section.
 4 Experiments
-4.1 Experimental Settings
-Table 1: Few-slice quantitative evaluation on unified dataset.
-We compared the predicted 3D meshes by deforming from ED frame to ES frame to the ground-truth 3D meshes.
-1- / 5-Slice: Number of slices used to recover the motion.
-Due to space limit, we omit the standard deviation.
                  1-Slice
@@
 Ours Tet 10.22 18.31 32.61 84.38 87.32 77.23 7.93 9.72 17.39 82.83 86.76 79.27 7.30 9.96 15.58 81.24 89.37 81.48
-Table 2: Clinical indexes evaluation on M&Ms dataset. 1- / 5- / Full-Slice denotes the number of slices used to recover motion.
+Table 1: Few-slice quantitative evaluation on unified dataset.
+We compared the predicted 3D meshes by deforming from ED frame to ES frame to the ground-truth 3D meshes.
+1- / 5-Slice: Number of slices used to recover the motion.
   Method LVESV MAE $\downarrow$ LVEF MAE $\downarrow$ RVESV MAE $\downarrow$ RVEF MAE $\downarrow$
@@
 Ours 4.84 2.04 9.65 4.77
-4.1.1 Data
+Table 2: Clinical indexes evaluation on M&Ms dataset. 1- / 5- / Full-Slice denotes the number of slices used to recover motion.
+4.1 Experimental Settings
+Datasets and Metrics.
 We train and validate our method on a unified dataset constructed from three large and diverse publicly available datasets: ACDC, M&Ms, and M&Ms-2 [5, 9, 25].
-ACDC features CMR sequences of about 30 frames for 100 patients, each one covering at least one cardiac cycle and with a slice thickness of 5-10 mm.
-For these datasets, segmentation masks for the left ventricle (LV), right ventricle (RV), and myocardium (MYO) are provided, but only for the end-diastolic (ED) and end-systolic (ES) frames.
-M&Ms and M&Ms-2 are made of multi-center and multi-vendor collected CMR sequences for 375 and 360 subjects, respectively.
-While featuring similar resolutions and sequence lengths as ACDC dataset, these two datasets cover a broader range of cardiac pathologies.
-As in earlier studies [13, 14, 9, 3], we register and fit a Statistic Shape Model (SSM) [24, 5] to segmentation mask at the ED and ES frames using non-rigid image registration methods to generate cardiac meshes.
-The unified dataset comprises 835 CMR sequences, randomly split into 501 (60%), 167 (20%), and 167 (20%) sequences for training, validation, and testing, respectively.
+Segmentation for the left ventricle (LV), right ventricle (RV), and myocardium (MYO) are provided, but only for the ED and ES frames.
+The unified dataset comprises 835 CMR sequences, randomly split into training, validation, and testing sets in a ratio of 60%, 20%, and 20%.
+Dataset details are provided in Appendix A.1.
 We use two additional datasets for external evaluation to assess generalization in a clinical context where model retraining is not feasible.
-The first dataset was collected at the Center for Interventional MRI, University Hospital Lausanne (CHUV), designed with the target online scenario in mind and hereafter referred to as CHUV.
-For each subject, a stack of short-axis (SAX) breath-hold cine slices was acquired at rest on a 1.5T interventional system (Siemens Healthineers, Magnetom Aera) using an accelerated steady-state free precession (SSFP) sequence covering the entire left ventricle.
-In addition, single SAX cine slices were acquired at mid-ventricular levels with slight angulations relative to the full SAX stack.
-Because both heart shape and rate may vary during intervention, updating of the 3D heart model is often required to ensure precise intracardiac manipulations.
-To mimic heart-rate changes, five volunteers exercised inside the interventional MRI (bending and extending their legs in two parallel plastic gutters for 2–3 minutes), thereby increasing their heart rates by approximately 20–30 beats per minute.
-Once elevated, single SAX cine slices were reacquired at the same mid-ventricular levels as during rest.
-For real-time acquisition, the online sequences had lower spatial resolution than the offline short-axis CMR stacks.
-To evaluate full-sequence performance across the entire cardiac cycle beyond the keyframes, we further use the second external dataset, 4DM [14], which focuses on left ventricular myocardium reconstruction and provides mesh annotations for every frame.
-This dataset enables a rigorous evaluation of the proposed weakly supervised motion learning scheme.
-4.1.2 Evaluation Metrics
-For evaluation, we compare the predicted 3D mesh obtained by deforming the ED frame to the ES frame with the ground-truth 3D mesh at the ES frame.
-We use L2-norm Chamfer Distance (CD) to measure 3D reconstruction quality.
-Dice score is also reported to quantify the intra-slice segmentation results.
-Besides these quality measures, we also report the Mean-Absolute Error (MAE) of LV/RV ES Volume (LVESV/RVESV), LV/RV ejection fraction (LVEF/RVEF) derived from our predicted meshes.
-LVEF and RVEF are representative cardiac function indexes which could be used for clinical diagnosis.
-4.1.3 Implementation Details
+The first dataset, referred to as iMRI, was collected on an interventional MRI system under both rest and elevated-heart-rate conditions.
+It is designed with the target online scenario in mind.
+For each subject, full short-axis cine stacks were acquired at rest, while additional single mid-ventricular slices were collected at rest and during mild in-scanner exercise to simulate heart-rate variations.
+The dataset thus reflects realistic cardiac motion dynamics encountered in online interventional MRI.
+Details on the scanning protocol, equipment, and exercise procedure are provided in Appendix A.2.
+To evaluate full-sequence performance across the entire cardiac cycle beyond the keyframes, we use the second external dataset, 4DM [52], which focuses on left ventricular myocardium reconstruction and provides mesh annotations for every frame.
+This dataset allows rigorous evaluation of the proposed weakly supervised motion learning scheme.
+For evaluation, we use the L2-norm Chamfer Distance (CD) to measure 3D reconstruction quality and the Dice score to quantify intra-slice segmentation results.
+We also report the Mean-Absolute Error of LV/RV ES Volume (LVESV/RVESV), LV/RV ejection fraction (LVEF/RVEF) derived from our predicted meshes.
+LVEF and RVEF are representative cardiac function indexes used for clinical diagnosis.
+See Appendix A.3 for details.
+Implementation Details.
 In all experiments, the slices are resampled by linear interpolation to a spacing of $1.25\times 1.25$ mm.
 From this, the AFA module constructs 3D volume with an effective spacing of $1.25\times 1.25\times 2$ mm.
-The AFA module uses a 8-head multi-head attention with learnable 3D position embedding.
-In the reconstruction branch, a modified nnU-Net [42] is used as the image encoder.
-It comprises five downsampling and upsampling blocks with channel sizes [32, 64, 128, 256, 320].
-We use the same GCN structure as in [18] with $3$ layers and $128$ channels.
-The initial tetrahedra resolution is $128$.
-For the motion branch, we use the identical encoding network structure as the reconstruction branch.
-For the deformation model, the hidden dimension of the GCN and GRU are both set to $128$.
-We use SGD optimizer with an initial learning rate of 0.01, a weight decay of 3e53e-5, and momentum of $0.99$.
-We train all models for $300/150$ epochs for the reconstruction/motion learning stage.
-All experiments are conducted using one V100 GPU.
-4.1.4 Baselines
+See Appendix B for model and optimization details.
+Baselines.
 We compare our method against several existing approaches for 4D shape reconstruction.
-For the online interventional setting, as noted earlier, most methods assume access to a complete CMR stack, making them unsuitable for the online few-slice setting.
-Therefore, we adapted several baselines so that they can operate under sparse inputs.
-4DMR [14] is an implicit representation method that reconstructs myocardium mesh sequence by optimizing decoupled shape and motion latent codes at test time, minimizing the discrepancy between the predicted mesh and ground-truth contour points.
-We adapt it by restricting the optimization to incomplete contours extracted from the available few slices, while still using full contours for the initial frame.
-MR-Net [26] is a template-mesh-based approach that takes ground-truth contour points as input and leverages a hybrid PointNet + 3D CNN architecture to predict vertex deformations.
-Since MR-Net directly encodes point sets, it can naturally accommodate incomplete inputs, a scenario already considered in the original paper.
-Although originally designed for static reconstruction, we adapt it for motion estimation by setting the ED-frame mesh as the template.
+For the online interventional setting, as noted earlier, most methods assume access to a complete CMR stack, making them unsuitable for few-slice inputs.
+Therefore, we adapted several baselines to operate under sparse inputs, including modified versions of 4DMR [52] and MR-Net [10] We also include two variants derived from shape reconstruction methods: MeshDeformNet [21] and DeepCSR [12].
+We denote them as Ours-Mesh and Ours-SDF since their main differences with ours lie in the shape representation.
+For the traditional offline setting, in addition to the above methods, we further compare against Free Form Deformation (FFD[38], diffeomorphic Demons (dDemons[44], MR-Net w.
+nnU-Net, MulViMotion and DeepMesh.
+Details about the modified baselines are provided in Appendix C.
 Figure 3: 1-Slice Comparison.
-(Left) We predict 3D meshes by deforming from ED to ES frame.
-Color indicates the magnitude of point-to-surface error in $m$.
+(Left) We predict meshes by deforming from ED to ES frame.
+Color indicates the magnitude of point-to-surface error.
 (Right) Segmentation results on slices at the apex and base location.
-Note that 4DMR can only output motion for myocardium.
-Figure 3: Motion sequence prediction results using different number of slices by deforming from ED frame to the target frame. Corresponding Volume-frame curve is given on the right. #1: NOR, #2: DCM.
-In addition, we include two modified baselines derived from shape reconstruction methods: MeshDeformNet [44] and DeepCSR [46].
-To adapt them for motion reconstruction, we integrated into both the same nnU-Net backbone, AFA module, deformation model, and training pipeline as our method.
-Since their main differences with our full framework lie in the shape representation, we denote them as Ours-Mesh and Ours-SDF.
-For the traditional offline setting, in addition to the above methods, we further compare against B-spline Free Form Deformation (FFD) [48], diffeomorphic Demons (dDemons) [49], MR-Net w.
-nnU-Net, MulViMotion and DeepMesh.
-FFD and dDemons are classical registration algorithms that have been widely used in recent cardiac motion tracking studies [50, 51, 52].
-MR-Net w.
-nnU-Net adopts the MR-Net architecture but replaces ground-truth contours with those predicted by nnU-Net, featuring a two-stage automated pipeline.
-MulViMotion and DeepMesh are mesh-based approaches that predict deformation fields via 3D CNNs, but only for the myocardium class.
-For fair comparison, we adopt their SAX-view-only versions as mentioned in the original papers.
-Unless explicitly stated otherwise, all methods are jointly trained on the unified dataset.
+Note 4DMR can only output motion for myocardium.
+Figure 4: Motion sequence prediction results using 1-/Full-slices by deforming from ED frame to the target frame. Corresponding Volume-frame curve is given on the right. #1: NOR, #2: DCM.
 4.2 Motion Estimation from a Few Slices
-Table 3: Full-slice quantitative evaluation on unified dataset. We compared the predicted 3D meshes by deforming from ED frame to ES frame to the ground-truth 3D meshes.
+We present results under the online few-slice scenario that represents the primary motivation of this work.
+Fig.
+3 shows a representative motion reconstruction using only the central slice as input: one from a normal (NOR) subject and another from a case of dilated cardiomyopathy (DCM).
+Due to stronger contractile function, the normal heart exhibits larger deformations, making motion prediction inherently more challenging, as reflected by higher reconstruction errors.
+Nevertheless, our method delivers accurate predictions in both cases.
+Moreover, segmentation results on slices near the apex and base demonstrate that our model produces reliable deformations even in regions distant from the input slice.
+Complete reconstructed sequences and their corresponding volume–time curves are shown in Fig.
+4.
+Quantitative results on the full test set are reported in Tab.
+1 and the top section of Tab.
+2, including Chamfer Distance, Dice scores, and cardiac function indices.
+For consistency, all experiments were performed using the central 1 or 5 slices, though slices at arbitrary positions could have been used, as discussed in Section 4.4.
+TetHeart consistently outperforms all baselines without requiring ground-truth contours or segmentations as input.
+It infers cardiac motion directly from raw CMR images, eliminating the need for time-consuming test-time optimization, and achieves an inference speed of 12 FPS on an NVIDIA V100 GPU.
+Further acceleration could be achieved through model quantization or lightweight encoders to meet strict real-time constraints.
+Notably, even when provided with a single slice, TetHeart produces meaningful physiological indices, underscoring its suitability for clinical deployment.
+Although Ours-Mesh and Ours-SDF incorporate the same sparse feature handling strategy as our full model, TetHeart still achieves superior performance.
+This confirms that the tetrahedral representation effectively preserves spatial coherence and serves as a more expressive basis for 3D motion inference under sparse observations.
+Fig.
+5 illustrates performance as a function of slice count: the fewer slices are used, the larger the gap between TetHeart and the ablated variants.
+Nevertheless, both variants still outperform all prior methods by a clear margin, demonstrating that the AFA module and the training strategy form a robust and generalizable solution for incomplete observations.
+The relatively poor performance of 4DMR can be attributed to its simple MLP architecture, which may lack the representational power to generalize well to the large and heterogeneous datasets considered here.
+Moreover, its test-time optimization strategy does not generalize well to severely under-sampled inputs.
+The weaker performance of MR-Net likely arises from the lack of reconstruction features and its sole reliance on contour points for deformation prediction, which overlooks rich cine image information.
+Figure 5: Chamfer distance as a function of the number of slices used to capture the motion.
+4.3 Motion Estimation from Full Stacks
   Method GT-Free Optim-Free Repre. ACDC M&Ms M&Ms-2
@@
 Ours Tet 6.63 11.51 22.15 87.12 88.96 84.21 4.11 4.85 13.68 84.31 87.43 83.99 4.02 5.04 10.82 85.33 89.13 86.94
-Table 4: Zero-shot evaluation results on public 4DM dataset. Check marks denote the dataset used for training. Oracle is trained and evaluated on 4DM.
+Table 3: Full-slice quantitative evaluation on unified dataset. We compared the predicted 3D meshes by deforming from ED frame to ES frame to the ground-truth 3D meshes.
+Since we had to modify existing techniques to compare them to ours in the few-slice scenario, we now compare our method to the original baselines on the full set of slices they were designed to operate on and report the results in Tab.
+3 and the bottom of Tab.
+2.
+In Fig.
+4 and S2, we show qualitative results on the same hearts as in Fig.
+3, but now using all slices as input instead of only one.
+Remarkably, although the main driver of this work was operating on as few slices as possible, we also consistently outperform the baselines in the full-slice setting.
+We attribute this to several factors.
+Using predicted contours in MR-Net leads to slightly worse performance, indicating such a two-stage automated pipeline is prone to error accumulation, consistent with the conclusions in the original paper [10].
+MulViMotion estimates voxel-wise motion fields and reconstructs meshes from warped segmentation maps, which is inherently less accurate than directly modeling meshes.
+Compared to DeepMesh, our model benefits from a shared encoding network, allowing the deformation branch to leverage features from the reconstruction branch.
+Moreover, the AFA module dynamically aggregates spatially adjacent features, improving prediction quality.
+4.4 Generalization without Retraining
+We now perform external evaluation to test the generalizability of our model to the iCMR and 4DM datasets without retraining.
+In Fig.
+6, we show sequence predictions on iCMR dataset with the spatial position of the real-time captured 2D slice visualized.
+As can be seen, TetHeart reconstructs plausible heart motion across the cardiac cycle under both rest and exercise situation.
+Moreover, the predicted volume–time curves confirm the physiological plausibility of our motion reconstruction.
+In the exercise condition, the heart exhibits a shorter ES–ED recovery period and a reduced duration of each cardiac cycle, aligns well with expectation.
+Please refer to Appendix E for more details and the external evaluation results on the 4DM dataset.
+Figure 6: Motion sequence prediction on the iCMR dataset.
+(Left) Spatial configuration of the initial SAX stack and the online observations.
+(Middle) Cardiac motion sequences for each subject under rest and exercise conditions.
+(Right) Volume–time curves.
+4.5 Ablation Studies
+We conduct ablation studies to evaluate our design choices.
+First, we examine the AFA module.
+Simply stacking 2D features yields similar performance to AFA in the full-slice setting but fails to generalize when the number or positions of slices vary, highlighting the need for a flexible aggregation mechanism.
+Removing position encodings moderately degrades performance, confirming the importance of spatial information, while using only position encodings leads to a substantial drop—showing that both image features and spatial coordinates are essential for effective motion prediction.
+When the motion branch is trained without parameter sharing and randomly initialized, performance decreases significantly, demonstrating that our shared-parameter design effectively transfers 3D knowledge from reconstruction to motion modeling.
+Lastly, we ablate the impact of the distillation loss.
+Without it, the model performs slightly better in the full-slice setting but encodes less discriminative features from sparse 2D inputs, resulting in poorer few-slice performance.
+Combining the distillation loss with our aggressive sampling strategy consistently improves results across all scenarios, underscoring the importance of this training scheme for improving robustness and generalization.
+Additional ablation results are provided in Appendix F.
+  Method M&Ms Myocardium CD (mm2) $\downarrow$
+1-Slice 5-Slice Full-Slice
+Ours 9.76 7.93 4.11
+- AFA module - - 4.22
+- position embedding 11.23 9.04 4.67
+- image feature in query $Q$ 16.67 12.37 6.44
+- shared encoding network 20.66 11.63 6.43
+- distillation loss 10.34 8.12 4.03
+Table 4: Ablation study results on M&Ms dataset.
+5 Conclusion
+We have presented TetHeart, a unified framework for 4D cardiac shape and motion reconstruction from 2D CMR slices, operating robustly from full stacks down to a single real-time slice.
+For the first time, we show that accurate 3D cardiac motion can be inferred from sparse intra-operative CMR data, enabling real-time tracking and guidance during interventions.
+This has the potential to shift cardiac reconstruction from retrospective analysis to live clinical application, laying the foundation for patient-specific digital twins of the beating heart and paving the way for adaptive, predictive, and personalized image-guided procedures.
+As an immediate next step, we plan to deploy TetHeart in an interventional MRI suite at our collaborating hospital.
+Looking ahead, we aim to extend this framework to other dynamic organs, integrate multimodal imaging, and develop predictive digital twins that anticipate motion under both physiological and interventional conditions—advancing toward a new generation of real-time, personalized, image-guided medicine.
+Supplementary Material
+Appendix A Datasets and Metrics
+A.1 Unified Dataset
+ACDC features CMR sequences of about 30 frames for 100 patients, each one covering at least one cardiac cycle and with a slice thickness of 5-10 mm.
+M&Ms and M&Ms-2 are made of multi-center and multi-vendor collected CMR sequences for 375 and 360 subjects, respectively.
+While featuring similar resolutions and sequence lengths as ACDC dataset, these two datasets cover a broader range of cardiac pathologies.
+For these datasets, segmentation masks for the left ventricle (LV), right ventricle (RV), and myocardium (MYO) are provided, but only for the end-diastolic (ED) and end-systolic (ES) frames.
+As in earlier studies [28, 52, 48, 33], we register and fit a Statistic Shape Model (SSM) [15, 2] to segmentation mask at the ED and ES frames using non-rigid image registration methods to generate cardiac meshes.
+A.2 iCMR Dataset
+The iCMR dataset was collected at the Center for Interventional MRI of our collaborating university hospital.
+For each subject, a stack of short-axis (SAX) breath-hold cine slices was acquired at rest on a 1.5T interventional system (Siemens Healthineers, Magnetom Aera) using an accelerated steady-state free precession (SSFP) sequence covering the entire left ventricle.
+In addition, single SAX cine slices were acquired at mid-ventricular levels with slight angulations relative to the full SAX stack.
+Because both heart shape and rate may vary during intervention, updating of the 3D heart model is often required to ensure precise intracardiac manipulations.
+To mimic heart-rate changes, five volunteers exercised inside the interventional MRI (bending and extending their legs in two parallel plastic gutters for 2–3 minutes), thereby increasing their heart rates by approximately 20–30 beats per minute.
+Once elevated, single SAX cine slices were reacquired at the same mid-ventricular levels as during rest.
+For real-time acquisition, the online sequences had lower spatial resolution than the offline short-axis CMR stacks to meet the latency constraint.
+A.3 Evaluation Metrics
+For evaluation, we compare the predicted 3D mesh obtained by deforming the ED frame to the ES frame with the ground-truth 3D mesh at the ES frame.
+We use L2-norm Chamfer Distance (CD) to measure 3D reconstruction quality, which is calculated as:
+$\mathrm{CD}(P,Q)=\frac{1}{|P|}\sum_{p\in P}\min_{q\in Q}\|p-q\|_{2}^{2}+\frac{1}{|Q|}\sum_{q\in Q}\min_{p\in P}\|q-p\|_{2}^{2}.$
+Dice score is also reported to quantify the intra-slice segmentation results, calculated as:
+$\mathrm{DSC}(A,B)=\frac{2|A\cap B|}{|A|+|B|}.$
+Besides these quality measures, we also report the Mean-Absolute Error (MAE) of LV/RV ES Volume (LVESV/RVESV), LV/RV ejection fraction (LVEF/RVEF) derived from our predicted meshes.
+The ejection fraction is calculated as
+$XEF=\frac{XEDV-XESV}{XEDV},X\in\{LV,RV\},$
+here EDV denotes the volume of the ED meshes.
+For our method, we use the predicted ED meshes from the reconstruction branch for calculation.
+For baselines require ground-truth input meshes, we directly use the input meshes for calculation.
+The MAE of ES volume and ejection fraction is then calculated as
+$MAE_{esv}=|ESV_{p}-ESV_{gt}|,~MAE_{ef}=|EF_{p}-EF_{gt}|.$
+The ground-truth volume and ejection fraction are derived from the ground-truth meshes. LVEF and RVEF are representative cardiac function indexes which could be used for clinical diagnosis.
+Appendix B Implementation Details
+In all experiments, the slices are resampled by linear interpolation to a spacing of $1.25\times 1.25$ mm.
+From this, the AFA module constructs 3D volume with an effective spacing of $1.25\times 1.25\times 2$ mm.
+The AFA module uses a 8-head multi-head attention with learnable 3D position embedding.
+In the reconstruction branch, a modified nnU-Net [20] is used as the image encoder.
+It comprises five downsampling and upsampling blocks with channel sizes [32, 64, 128, 256, 320].
+We use the same GCN structure as in [40] with $3$ layers and $128$ channels.
+The initial tetrahedra resolution is $128$.
+For the motion branch, we use the identical encoding network structure as the reconstruction branch.
+For the deformation model, the hidden dimension of the GCN and GRU are both set to $128$.
+We use SGD optimizer with an initial learning rate of 0.01, a weight decay of 3e53e-5, and momentum of $0.99$.
+We train all models for $300/150$ epochs for the reconstruction/motion learning stage.
+All experiments are conducted using one V100 GPU.
+Figure S1: Motion sequence prediction results using 1-/Full-slices by deforming from ED frame to the target frame. Corresponding Volume-frame curve is given on the right. #1: NOR, #2: DCM.
+Appendix C Baselines
+We compare our method against several existing approaches for 4D shape reconstruction.
+For the online interventional setting, we adapted several baselines so that they can operate under sparse inputs.
+4DMR [52] is an implicit representation method that reconstructs myocardium mesh sequence by optimizing decoupled shape and motion latent codes at test time, minimizing the discrepancy between the predicted mesh and ground-truth contour points.
+We adapt it by restricting the optimization to incomplete contours extracted from the available few slices, while still using full contours for the initial frame.
+MR-Net [10] is a template-mesh-based approach that takes ground-truth contour points as input and leverages a hybrid PointNet + 3D CNN architecture to predict vertex deformations.
+Since MR-Net directly encodes point sets, it can naturally accommodate incomplete inputs, a scenario already considered in the original paper.
+Although originally designed for static reconstruction, we adapt it for motion estimation by setting the ED-frame mesh as the template.
+In addition, we include two modified baselines derived from shape reconstruction methods: MeshDeformNet [21] and DeepCSR [12].
+To adapt them for motion reconstruction, we integrated into both the same nnU-Net backbone, AFA module, deformation model, and training pipeline as our method.
+Since their main differences with our full framework lie in the shape representation, we denote them as Ours-Mesh and Ours-SDF.
+For the traditional offline setting, in addition to the above methods, we further compare against B-spline Free Form Deformation (FFD) [38], diffeomorphic Demons (dDemons) [44], MR-Net w.
+nnU-Net, MulViMotion and DeepMesh.
+FFD and dDemons are classical registration algorithms that have been widely used in recent cardiac motion tracking studies [3, 35, 31].
+MR-Net w.
+nnU-Net adopts the MR-Net architecture but replaces ground-truth contours with those predicted by nnU-Net, featuring a two-stage automated pipeline.
+MulViMotion and DeepMesh are mesh-based approaches that predict deformation fields via 3D CNNs, but only for the myocardium class.
+For fair comparison, we adopt their SAX-view-only versions as mentioned in the original papers.
+Unless explicitly stated otherwise, all methods are jointly trained on the unified dataset.
+Figure S2: Full-Slice Comparison.
+(Left) We predict 3D meshes by deforming from ED to ES frame.
+Color indicates the magnitude of point-to-surface error in $m$.
+(Right) Segmentation results on slices at the apex and base location.
+Note 4DMR and DeepMesh can only output for myocardium.
   Method Training Dataset Myocardium CD (mm2) $\downarrow$
@@
 Ours (Oracle) - - - 6.24 4.87 2.70
-In this section we present results under the online few-slice scenario that has been the main motivation for this work.
-In Fig.
-2, we show two representative motion reconstruction results using only the central slice as input: one from a normal (NOR) heart and another from a subject with dilated cardiomyopathy (DCM).
-Because of its stronger contractile function, the normal heart exhibits larger deformations than the DCM heart, making motion prediction more challenging, which is reflected by larger magnitudes in the error maps.
-Nevertheless, our method delivers accurate predictions in both cases.
-Furthermore, the segmentation results on slices at the apex and base location demonstrate that our model can provide reliable deformation estimation even in regions far from the slices used for motion inference.
-The complete reconstructed sequences together with the corresponding volume–time curves are shown in Fig.
-3.
-We report quantitative results on the whole test set in Tabs.
-1 and at the top of Tab.
-2.
-The first contains chamfer distances and dice scores and the second cardiac function index predictions.
-For consistency, we generated all these results using the central 1 or 5 slices even though slices at any position could be used, as shown in Section 4.4.
-TetHeart consistently outperforms our four baselines without requiring ground-truth segmentations or contours as input.
-Instead, it infers cardiac motion directly from raw CMR images without any time-consuming test-time optimization.
-As a result, it delivers fast inference—12 FPS on a single V100—with a potential for further acceleration by adopting a lightweight encoder architecture and leveraging model quantization techniques to meet the real-time constraints.
-Even when using a single slice, TetHeart delivers meaningful indices.
-All of this points toward its suitability for eventual clinical deployment.
-Figure 4: Chamfer distance as a function of the number of slices used to capture the motion.
-Although we incorporated into Ours-Mesh and Ours-SDF the same mechanisms for handling sparse features as in our complete method, the latter still delivers better results.
-This confirms that the tetrahedral representation we use is more effective at preserving spatial information within the object, making it a better choice for inferring 3D motion when only limited observations are available.
-To quantify this, we plot in Fig.
-4 the performance of all three methods as a function of the number of slices used.
-The fewer, the greater the gap between our complete approach and the other two.
-Nevertheless, both variants still surpass prior approaches by a clear margin, demonstrating that design components such as the AFA module and our training strategy form a general and effective solution for handling incomplete slice scenarios.
-The comparatively poor performance of 4DMR can be attributed to its reliance on a simple MLP architecture, which may lack the representational power to generalize well to the large and heterogeneous datasets considered in this study.
-Moreover, its test-time optimization strategy does not generalize well to severely under-sampled inputs.
-The weaker performance of MR-Net likely stems from the fact that it reconstructs meshes solely from contour points, thereby neglecting the rich information available in cine images.
-4.3 Motion Estimation from Full Stacks
-Figure 5: Full-Slice Comparison.
-(Left) We predict 3D meshes by deforming from ED to ES frame.
-Color indicates the magnitude of point-to-surface error in $m$.
-(Right) Segmentation results on slices at the apex and base location.
-Note 4DMR and DeepMesh can only output for myocardium.
-Since we had to modify existing techniques to compare them to ours in the few-slice scenario, we now compare our method to the original baselines on the full set of slices they were designed to operate on and report the results in Tab.
-3 and the bottom of Tab.
-2.
-In Fig.
-3 and 5, we show qualitative results on the same hearts as in Fig.
-2, but now using all slices as input instead of only one.
-Remarkably, even though the main driver of this work was operating on as few slices as possible, we also consistently outperform the baselines in the full-slice scenario.
-We attribute this to several factors.
-Consistent with the previous analysis, the relatively poor performance of MR-Net may be attributed to its neglect of the rich information available in cine images.
-Using predicted contours resulted in slightly inferior performance, indicating that such a two-stage automated pipeline is prone to error accumulation, a finding consistent with the conclusions in the original paper [26].
-MulViMotion estimates voxel-wise motion fields and reconstructs meshes from warped segmentation maps, which is inherently less accurate than directly modeling meshes.
-Compared to DeepMesh, our model benefits from a shared encoding network, allowing the deformation branch to leverage informative features from the reconstruction branch.
-Moreover, our AFA module dynamically aggregates spatially adjacent features, further enhancing prediction quality.
-4.4 External Evaluation: To Generalize without Retraining
-We now test the generalizability of our model to the CHUV and 4DM datasets without retraining.
-In Fig.
-6, we show sequence predictions on CHUV with the spatial position of the real-time captured 2D slice visualized in the left.
-As can be seen, TetHeart reconstructs a plausible heart motion across the cardiac cycle under both rest and exercise situation.
-Moreover, the predicted volume–time curves further confirm the physiological plausibility of our motion reconstruction.
-In the exercise condition, the heart exhibits a shorter ES–ED recovery period and a reduced duration of each cardiac cycle, which aligns well with expectation.
+Table S1: Zero-shot evaluation results on public 4DM dataset. Check marks denote the dataset used for training. Oracle is trained and evaluated on 4DM.
+Appendix D Complementary Experiment Results
+High-Res Verision of Fig. 4. In Fig. S1, we provide a higher resolution version of Fig. 4 for better view.
+Disease Definition.
+A normal heart should satisfy the following criteria: LV EF $50\%$ and RV EF $40\%$.
+A subject with dilated cardiomyopathy should have LV EF $40\%$.
+This is not a strict definition; a more rigorous classification criterion would require a comprehensive analysis of the ratio between volume and body surface area, along with other indicators.
+However, verifying whether the predicted ejection fraction from TetHeart satisfies this approximate criterion can further demonstrate the practicality of our model.
+As could be seen from Fig. S1, even with a single slice, TetHeart could predict cardiac function indexes that meet the criteria. This highlights our model’s practicality.
+Visualization Results Using Full Slices. In Fig. S2, we show qualitative results on the same hearts as in Fig. 3. But now using full slices rather than using a single slice.
+Appendix E External Evaluation Results
+iCMR Dataset. We provide a higher resolution version of Fig. 6 in Fig. S3 for better view.
+Figure S3: Motion sequence prediction on the iCMR dataset.
+(Left) Spatial configuration of the initial SAX stack and the online observations.
+(Middle) Cardiac motion sequences for each subject under rest and exercise conditions.
+(Right) Volume–time curves.
+4DM Dataset.
 As 4DM provides full-sequence annotations, but only for the myocardium.
 we evaluate the full-sequence prediction performance for the myocardium subclass, rather than focusing solely on the ES frame and report the results in Tab.
-4.
+S1.
 Here, check marks denote the dataset used for training.
@@
 In contrast, 4DMR exhibits poor generalization, likely due to significant distribution shifts across datasets and the limited representational capacity of its simple MLP architecture.
-While MulViMotion and DeepMesh show some generalization ability, their lower baseline performance causes them to lag behind our method.
+While MulViMotion and DeepMesh show some generalization ability, their lower baseline performance causes them to lag behind us.
 Moreover, training with a larger and more diverse dataset improves generalization.
-And consistent with the findings from previous evaluations, representing the heart with tetrahedra enables more effective spatial information retention and exchange, particularly in scenarios with limited slice input.
-Figure 6: Motion sequence prediction on the CHUV dataset.
-(Left) Spatial configuration of the initial SAX stack and the online observations.
-(Middle) Cardiac motion sequences for each subject under rest and exercise conditions.
-(Right) Volume–time curves.
-4.5 Ablation Studies
-Finally, we conduct ablation studies to evaluate our design choices.
-First, we examine the role of the AFA module.
-Simply stacking 2D features achieves comparable performance to AFA in the full-slice setting, but it fails to generalize to scenarios with variable numbers and positions of 2D slices.
-This highlights the necessity of a more flexible mechanism like AFA.
-Next, we evaluate the importance of spatial information in the AFA module.
-If spatial position encodings are removed from the attention computation, the model’s performance degrades moderately, indicating that spatial information is crucial for accurate feature aggregation.
-However, if we use only position encodings as the query and discard image features entirely, the performance drops substantially.
-This shows that both image features and spatial coordinates are essential and must work together for effective motion prediction.
-When training the motion branch, if we do not share parameters with the reconstruction branch and instead initialize the deformation branch randomly, the model performance drops significantly.
-This demonstrates the effectiveness of our parameter-sharing strategy: the 3D knowledge learned during reconstruction is successfully transferred to motion modeling.
-Lastly, we ablate the impact of the distillation loss.
-Without it, the model achieves slightly better results in the full-slice setting, but the features encoded from sparse 2D inputs are less discriminative, leading to poorer performance under few-slice conditions.
-When the distillation loss is combined with our aggressive sampling strategy, the model achieves consistently stronger results across all settings, demonstrating the importance of this training scheme for improving robustness and generalization.
-Table 5: Ablation study results on M&Ms dataset.
-  Method M&Ms Myocardium CD (mm2) $\downarrow$
+And consistent with the findings from previous evaluations, representing the heart as tetrahedra enables more effective spatial information retention and exchange, particularly in scenarios with limited slice input.
+Appendix F Extended Ablation Studies
+The Impact of AFA Configuration.
+In Tab.
+S2, we show the impact of the number of slices in the AFA module and the number of selected positions per slice on both performance and speed (measured on an NVIDIA V100 GPU).
+As shown in the table, our default configuration achieves a good balance between accuracy and efficiency.
+Further increasing the number of slices or selected positions brings only limited performance gains but leads to a slowdown.
+One thing worth mentioning is that to implement our AFA module we use a relatively simple strategy of sorting and selecting the top-k elements.
+Therefore, for $|S|=1,|G|=5^{2}$ and $|S|=3,|G|=3^{2}$, although the two configurations use a similar number of features, the latter runs slower because it performs sorting multiple times.
+More advanced strategies or precomputation could potentially narrow this gap for further acceleration.
+  Model $|S|$ $|G|$ Myocardium CD & FPS ($mm^{2}$ $\downarrow$ & $img/s$ $\uparrow$)
 1-Slice 5-Slice Full-Slice
-Ours 9.76 7.93 4.11
-- AFA module - - 4.22
-- position embedding 11.23 9.04 4.67
-- image feature in query $Q$ 16.67 12.37 6.44
-- shared encoding network 20.66 11.63 6.43
-- distillation loss 10.34 8.12 4.03
-5 Conclusion
-We have proposed TetHeart, a unified framework for reconstructing heart motion from either full stacks of 2D MRI slices or from sparse sets, down to a single slice.
-We demonstrate, for the first time, that full 3D heart shape and motion can be inferred in sparse real-time CMR slices, making our approach suitable for guidance of interventional procedures.
-This breakthrough shifts cardiac reconstruction from retrospective analysis to live clinical use, laying the foundation for real-time digital twins of the beating heart.
-This capability not only has the potential to enable precise, image-guided interventions today, but also points to a future where digital twins of organs in motion guide therapies across medicine—adaptive, predictive, and personalized at the point of care.
-[…]
+  TetHeart 1 $1^{2}$ 10.34 15.1 8.45 15.1 4.56 15.1
+1 $3^{2}$ 9.98 15.0 8.22 14.9 4.43 14.9
+1 $5^{2}$ 9.93 14.8 8.13 14.7 4.39 14.6
+3 $1^{2}$ 10.32 15.0 8.19 13.0 4.37 13.0
+3* $3^{2}*$ 9.76 14.9 7.93 12.1 4.11 12.0
+3 $5^{2}$ 9.73 14.7 7.86 10.8 4.05 10.7
+5 $1^{2}$ 10.27 15.0 8.17 11.9 4.33 11.8
+5 $3^{2}$ 9.74 14.8 7.81 10.4 4.09 10.4
+5 $5^{2}$ 9.72 14.7 7.83 8.7 4.10 8.7
+7 $5^{2}$ 9.76 14.6 7.85 8.6 4.07 7.3
+Table S2: Impact of different AFA configuration evaluated on the M&Ms dataset. $|S|$: the number of slice. $|G|$: the number of positions selected from each slice. *: default setting.
+The Influence of the Input Slice Position.
+In the main text, for ease of comparison and to ensure fairness, we use central slices as model inputs to evaluate the performance of different methods.
+Here we investigate how the input slice position affects the accuracy of motion inference.
+The results are shown in Tab.
+S3.
+The numbers indicate the offset of the input slice relative to the central slice: - denotes a shift toward the apex, and ++ denotes a shift toward the base.
+For example, suppose we have total 11 slices, then the central slice is the 6-th slice, the +1+1 slice is the 7-th slice, the 1-1 slice is the 5-th slice.
+As seen from the table, although the model still achieves good motion inference performance, its accuracy steadily decreases as the offset increases.
+These experimental results suggest that slices at different positions contain different amounts of motion information, which may be more focused on local regions.
+If only single slice is available, selecting it close to the central position yields better overall motion inference.
+This conclusion provides guidance on choosing suitable scan positions in real-world intervention workflows.
+For multiple slices, we could conduct similar experiments to explore which slice combinations are most effective.
+However, due to the large number of possible combinations, we provide only one general conclusion here: using an appropriate slice sampling interval can improve model performance compared with using adjacent slices.
+  Slice Position -3 -2 -1 Central +1 +2 +3
+  Myo CD ($mm^{2}$ $\downarrow$) 15.76 13.62 10.01 9.76 10.29 10.81 12.22
+Table S3: Impact of input slice position evaluated on the M&Ms dataset.
+The numbers indicate the offset of the input slice relative to the central slice.
+- denotes a shift towards the apex, and ++ denotes a shift towards the base.
+Category-Wise Ablation Results.
+In Tab.
+S4, we present more detailed results on the unified dataset as a supplement to Tab.
+4 in the main text.
+As shown in the table, across all datasets and all categories, the conclusions remain consistent with those in the main text: the design of the AFA module, parameter sharing, and the use of distillation loss all contribute to improving the final performance of our model.
+  Method Myocardium CD (mm2) $\downarrow$ LV CD (mm2) $\downarrow$ RV CD (mm2) $\downarrow$
+1-Slice 5-Slice Full-Slice 1-Slice 5-Slice Full-Slice 1-Slice 5-Slice Full-Slice
+ACDC
+Ours 15.24 10.22 6.63 23.65 18.31 11.51 40.64 32.61 22.15
+- AFA module - - 6.55 - - 11.58 - - 22.30
+- position embedding 16.60 11.33 6.84 25.32 20.14 12.73 42.23 34.19 23.59
+- image feature in query $Q$ 24.10 15.96 9.39 36.76 28.39 17.44 58.31 43.21 32.32
+- shared encoding network 32.23 15.24 10.01 49.12 28.99 18.59 75.92 44.81 34.44
+- distillation loss 16.13 10.48 6.49 25.07 18.78 11.28 43.08 33.43 21.71
+M&Ms
+Ours 9.76 7.93 4.11 12.37 9.72 4.85 21.42 17.39 13.68
+- AFA module - - 4.22 - - 4.79 - - 13.74
+- position embedding 11.23 9.04 4.67 13.35 10.34 5.67 22.72 18.84 14.93
+- image feature in query $Q$ 16.67 12.37 6.44 19.76 14.17 7.82 33.63 23.81 20.60
+- shared encoding network 20.66 11.63 6.43 24.56 14.24 7.77 41.80 24.12 20.45
+- distillation loss 10.34 8.12 4.03 13.11 9.94 4.75 22.70 17.79 13.41
+M&Ms-2
+Ours 10.12 7.30 4.02 13.89 9.96 5.04 20.27 15.58 10.82
+- AFA module - - 4.17 - - 5.07 - - 10.76
+- position embedding 11.93 8.41 4.74 14.98 10.40 5.69 21.91 16.72 11.65
+- image feature in query $Q$ 17.34 11.38 6.31 22.32 13.57 7.57 32.65 21.40 15.49
+- shared encoding network 21.49 11.09 6.54 26.96 13.64 7.85 39.44 20.73 16.07
+- distillation loss 10.95 7.47 4.11 15.01 10.19 5.05 21.89 15.94 11.06
+Table S4: More comprehensive ablation study results on the unified dataset.
+Appendix G Limitation
+Our method achieve promising results on both full-stack and few-slice settings, surpassing previous methods.
+However, by examining our own model, we can observe that although we have proposed several techniques, there still remains a performance gap between motion reconstruction using only a few slices versus the full stack.
+We plan to introduce physical constraints as prior knowledge into the model to reduce this gap.
+Another choice is to use a generative model to generate the mesh sequence progressively to complement our current feed-forward reconstruction process.
+However we need to strike a good balance between the reconstruction speed and quality as we are targeting real-time applications.
+Although we report quantitative results for the few-slice setting on the Unified dataset and achieve the best performance, we can only provide qualitative dynamic reconstruction results on the iCMR dataset because the ground-truth cardiac shapes for single slices are not available.
+Identifying an effective evaluation metric for assessing the quality of reconstructed mesh sequences in this setting would be very helpful for future clinical deployment.
+These are all interesting research directions, and we leave them for future work.